### v1

# Deblurring 3D Gaussian Splatting

Abstract
Recent studies in Radiance Fields have paved the robust way for novel view synthesis with their photorealistic rendering quality. Nevertheless, they usually employ neural networks and volumetric rendering, which are costly to train and impede their broad use in various real-time applications due to the lengthy rendering time. Lately 3D Gaussians splatting-based approach has been proposed to model the 3D scene, and it achieves remarkable visual quality while rendering the images in real-time. However, it suffers from severe degradation in the rendering quality if the training images are blurry. Blurriness commonly occurs due to the lens defocusing, object motion, and camera shake, and it inevitably intervenes in clean image acquisition. Several previous studies have attempted to render clean and sharp images from blurry input images using neural fields. The majority of those works, however, are designed only for volumetric rendering-based neural radiance fields and are not straightforwardly applicable to rasterization-based 3D Gaussian splatting methods. Thus, we propose a novel real-time deblurring framework, deblurring 3D Gaussian Splatting, using a small Multi-Layer Perceptron (MLP) that manipulates the covariance of each 3D Gaussian to model the scene blurriness. While deblurring 3D Gaussian Splatting can still enjoy real-time rendering, it can reconstruct fine and sharp details from blurry images. A variety of experiments have been conducted on the benchmark, and the results have revealed the effectiveness of our approach for deblurring. Qualitative results are available at https://benhenryl.github.io/Deblurring-3D-Gaussian-Splatting/
1 Introduction
With the emergence of Neural Radiance Fields (NeRF) [23], Novel view synthesis (NVS) has accounted for more roles in computer vision and graphics with its photorealistic scene reconstruction and applicability to diverse domains such as augmented/virtual reality(AR/VR) and robotics. Various NVS methods typically involve modeling 3D scenes from multiple 2D images from arbitrary viewpoints, and these images are often taken under diverse conditions. One of the significant challenges, particularly in practical scenarios, is the common occurrence of blurring effects. It has been a major bottleneck in rendering clean and high-fidelity novel view images, as it requires accurately reconstructing the 3D scene from the blurred input images.
NeRF [23] has shown outstanding performance in synthesizing photo-realistic images for novel viewpoints by representing 3D scenes with implicit functions. The volume rendering [7] technique has been a critical component of the massive success of NeRF. This can be attributed to its continuous nature and differentiability, making it well-suited to today’s prevalent automatic differentiation software ecosystems. However, significant rendering and training costs are associated with the volumetric rendering approach due to its reliance on dense sampling along the ray to generate a pixel, which requires substantial computational resources. Despite the recent advancements [10, 38, 24, 8, 9] that significantly reduce training time from days to minutes, improving the rendering time still remains a vital challenge.
Figure 1: Performance comparison to state-of-the-art deblurring NeRFs. Ours achieved a fast rendering speed (200 fps vs. 1 fps) while maintaining competitive rendered image quality (x-axis is represented in log scale).
Recently, 3D Gaussian Splatting (3D-GS) [15] has gained significant attention, demonstrating a capability to produce high-quality images at a remarkably fast rendering speed. It combines a large number of colored 3D Gaussians to represent 3D scenes with a differentiable splatting-based rasterization, substituting NeRF’s time-demanding volumetric rendering. The rendering process projects 3D Gaussian points onto the 2D image planes, and the positions, sizes, rotations, colors, and opacities of these Gaussians are adjusted via gradient-based optimization to better capture the 3D scenes. 3D-GS leverages rasterization, which can be significantly more efficient than volume rendering techniques on modern graphics hardware, thereby enabling rapid real-time rendering.
Expanding on the impressive capabilities of 3D-GS, we aim to further improve its robustness and versatility for more realistic settings, especially those involving blurring effects. Several approaches have attempted to handle the blurring issues in the recent NeRF literature [22, 20, 6, 41, 43]. The pioneering work is Deblur-NeRF [22], which renders sharp images from the neural radiance fields and uses an extra multi-layer perceptron (MLP) to produce the blur kernels. DP-NeRF [20] constrains neural radiance fields with two physical priors derived from the actual blurring process to reconstruct clean images. PDRF [28] uses a two-stage deblurring scheme and a voxel representation to further improve deblurring and training time. All works mentioned above have been developed under the assumption of volumetric rendering, which is not straightforwardly applicable to rasterization-based 3D-GS. Another line of works [41, 6], though not dependent on volume rendering, only address a specific type of blur, i.e., camera motion blur, and are not valid for mitigating the defocus blur.
In this work, we propose Deblurring 3D-GS, the first defocus deblurring algorithm for 3D-GS, which is well aligned with rasterization and thus enables real-time rendering. To do so, we modify the covariance matrices of 3D Gaussians to model the blurriness. Specifically, we employ a small MLP, which manipulates the covariance of each 3D Gaussian to model the scene blurriness. As blurriness is a phenomenon that is based on the intermingling of the neighboring pixels, our Deblurring 3D-GS simulates such an intermixing during the training time. To this end, we designed a framework that utilizes an MLP to learn the variations in different attributes of 3D Gaussians. These small variations are multiplied to the original values of the attributes, which in turn determine the updated shape of the resulting Gaussians. During the inference time, we render the scene using only the original components of 3D-GS without any additional offsets from MLP; thereby, 3D-GS can render sharp images because each pixel is free from the intermingling of nearby pixels. Further, since the MLP is not activated during the inference time, it can still enjoy real-time rendering similar to the 3D-GS while it can reconstruct fine and sharp details from the blurry images.
Figure 2: Our method’s overall workflow. $\gamma(\cdot)$ denotes positional encoding, $\odot$ denotes hadamard product, and $x$, $r$, $s$ stand for position, quaternion, and scaling of 3D Gaussian respectively. Given a set of 3D Gaussians $G$, we extract $x$, $r$, $s$ from each 3D Gaussian. Viewing direction $v$ from the dataset and $x$ are positionally encoded and then fed to $F_{\theta}$, a small MLP parameterized with $\theta$, which yields offsets $\delta r$ and $\delta s$ for each Gaussian. These $\delta r$ and $\delta s$ are element-wisely multiplied to their respective Gaussians’s $r$ and $s$ respectively, and these computed attributes are used to construct transformed 3D Gaussians $G^{\prime}(x,r\cdot\delta r,s\cdot\delta s)$. These transformed Gaussians $G^{\prime}$ are differentiably rasterized to render blurred image during training time. However, for inference stage, $G$ is directly fed to differentiable rasterizer, without involving any MLP, to render sharp image.
3D-GS [15] models a 3D scene from a sparse point cloud, which is usually obtained from the structure-from-motion (SfM) [35]. SfM extracts features from multi-view images and relates them via 3D points in the scene. If the given images are blurry, SfM fails heavily in identifying the valid features, and ends up extracting a very small number of points. Even worse, if the scene has a larger depth of field, SfM hardly extracts any points which lie on the far end of the scene. Due to this excessive sparsity in the point cloud constructed from set of blurry images, existing methods, including 3D-GS [15], that rely on point clouds fail to reconstruct the scene with fine details. To compensate for this excessive sparsity, we propose to add extra points with valid color features to the point cloud using N-nearest-neighbor interpolation [29]. In addition, we prune Gaussians based on their position to keep more Gaussians on the far plane.
A variety of experiments have been conducted on the benchmark, and the results have revealed the effectiveness of our approach for deblurring. Tested under different evaluation matrices, our method achieves state-of-the-art rendering quality or performs on par with the currently leading models while achieving significantly faster rendering speed ($>$ 200 FPS)
To sum up, our contributions are the following:
  • We propose the first real-time rendering-enabled defocus deblurring framework using 3D-GS.
  • We propose a novel technique that manipulates the covariance matrix of each 3D Gaussian differently to model spatially changing blur using a small MLP.
  • To compensate for sparse point clouds due to the blurry images, we propose a training technique that prunes and adds extra points with valid color features so that we can put more points on the far plane of the scene and harshly blurry regions.
  • We achieve FPS $>$ 200 while accomplishing superior rendering quality or performing on par with the existing cutting-edge models under different metrics.
2 Related Works
2.1 Image Deblurring
It is common to observe that when we casually take pictures with optical imaging systems, some parts of scene appear blurred in the images. This blurriness is caused by a variety of factors, including object motion, camera shake, and lens defocusing [1, 34]. When the image plane is separated from the ideal reference plane during the imaging process, out-of-focus or defocus blur happens. The degradation induced by defocus blur of an image is generally expressed as follows:
$g(x)=\sum_{s\in S_{h}}h(x,s)f(x)+n(x),~{}x\in S_{f},$ (1)
where $g(x)$ represents an observed blurry image, $h(x,s)$ is a blur kernel or Point Spread Function (PSF), $f(x)$ is a latent sharp image, and $n(x)$ denotes an additive white Gaussian noise that frequently occurs in nature images. $S_{f}\subset\mathbb{R}^{2}$ is a support set of an image and $S_{h}\subset\mathbb{R}^{2}$ is a support set of a blur kernel or PSF [18].
The first canonical approach to deconvolve an image is [33]. It applies iterative minimization to an energy function to obtain a maximum likelihood approximation of the original image, using a known PSF (the blur kernel). Typically, the blurring kernel $K$ and the sharp image $I$ are unknown. The technique termed as image blind deblurring [42] involves recovering the latent sharp image given just one blurry image $B$. In this procedure, the deblurred image is obtained by first estimating a blurring kernel. Image blind deblurring is a long-standing and ill-posed problem in the field of image and vision. Numerous approaches to address the blurring issue have been suggested in the literature. Among them, traditional methods often construct deblurring as an optimization problem and rely on natural image priors [21, 26, 44, 47]. Conversely, the majority of deep learning-based techniques use convolutional neural networks (CNN) to map the blurry image with the latent sharp image directly [25, 46, 31]. While series of studies have been actively conducted for image deblurring, they are mainly designed for deblurring 2D images and are not easily applicable to 3D scenes deblurring due to the lack of 3D view consistency.
2.2 Neural Radiance Fields
Neural Radiance Fields (NeRF) are a potent method that has recently gained popularity for creating high-fidelity 3D scenes from 2D images. Realistic rendering from novel viewpoints is made possible by NeRF, which uses deep neural networks to encode volumetric scene features. To estimate density $\sigma\in[0,\infty)$ and color value $c\in[0,1]^{3}$ of a given point, a radiance field is a continuous function $f$ that maps a 3D location $x\in\mathbb{R}^{3}$ and a viewing direction $d\in\mathbb{S}^{2}$. This function has been parameterized by a multi-layer perceptron (MLP) [23], where the weights of MLP are optimized to reconstruct a series of input photos of a particular scene: $(c,\sigma)=f_{\theta}:\big{(}\gamma(x),\gamma(d)\big{)}$ Here, $\theta$ indicates the network weights, and $\gamma$ is the specified positional encoding applied to $x$ and $d$ [39]. To generate the images at novel views, volume rendering [7] is used, taking into account the volume density and color of points.
2.2.1 Fast Inference NeRF
Numerous follow-up studies have been carried out to enhance NeRF’s rendering time to achieve real-time rendering. Many methods, such as grid-based approaches [40, 4, 38, 3, 8, 32], or those relying on hash [24, 2] adopt additional data structures to effectively reduce the size and number of layers of MLP and successfully improve the inference speed. However, they still fail to reach real-time view synthesis. Another line of works [30, 13, 45] proposes to bake the trained parameters into the faster representation and attain real-time rendering. While these methods rely on volumetric rendering, recently, 3D-GS [15] successfully renders photo-realistic images at novel views with noticeable rendering speed using a differentiable rasterizer and 3D Gaussians. Although several approaches have attempted to render tens or hundreds of images in a second, deblurring the blurry scene in real-time is not addressed, while blurriness commonly hinders clean image acquisition in the wild.
2.2.2 Deblurring NeRF
Several strategies have been proposed to train NeRF to render clean and sharp images from blurry input images. While DoF-NeRF [43] attempts to deblur the blurry scene, both all-in-focus and blurry images are required to train the model. Deblur-NeRF [22] firstly suggests deblurring NeRF without any all-in-focus images during training. It employs an additional small MLP, which predicts per-pixel blur kernel. However, it has to render the neighboring pixels of the target pixel in advance to model the interference of the neighboring information, so it takes additional computation at the time of training. Though the inference stage does not involve the blur kernel estimation, it is no different from the training with regard to rendering time as it is based on volumetric rendering which takes several seconds to render a single image. DP-NeRF [19] improved Deblur-NeRF in terms of image quality, and PDRF [28] partially replaces MLP to the grid and ameliorates rendering speed, still they depend on volumetric rendering to produce images and are not free from the rendering cost. Other works [6, 41] are aimed at addressing another type of blur, camera motion blur, rather than solving the long rendering time. While these deblurring NeRFs successfully produce clean images from the blurry input images, there is room for improvement in terms of rendering time. Thus, we propose a novel deblurring framework, deblurring 3D Gaussian Splatting, which enables real-time sharp image rendering using a differentiable rasterizer and 3D Gaussians.
3 Deblurring 3D Gaussian Splatting
Based on the 3D-GS [15], we generate 3D Gaussians, and each Gaussian is uniquely characterized by a set of the parameters, including 3D position $x$, opacity $\sigma$, and covariance matrix derived from quaternion $r$ scaling $s$. Each 3D Gaussian also contains spherical harmonics (SH) to represent view-dependent appearance. The input for the proposed method consists of camera poses and point clouds, which can be obtained through the structure from motion (SfM) [35], and a collection of images (possibly blurred). To deblur 3D Gaussians, we employ an MLP that takes the positions of the 3D Gaussians along with the $r$ and $s$ as inputs and outputs $\delta r$, and $\delta s$ which are the small scaling factors multiplied to $r$ and $s$, respectively. With new quaternion and scale, $r\cdot\delta r$ and $s\cdot\delta s$, the updated 3D Gaussians are subsequently fed to the tile-based rasterizer to produce the blurry images. The overview of our method is shown in Fig. 2.
3.1 Differential Rendering via 3D Gaussian Splatting
At the training time, the blurry images are rendered in a differentiable way and we use a gradient-based optimization to train our deblurring 3D Gaussians. We adopt methods from [15], which proposes differentiable rasterization. Each 3D Gaussian is defined by its covariance matrix $\Sigma(r,s)$ with mean value in 3D world space $x$ as following:
$G(x,r,s)=e^{-\frac{1}{2}x^{T}\Sigma^{-1}(r,s)x}.\vspace{0.2cm}$ (2)
Besides $\Sigma(r,s)$ and $x$, 3D Gaussians are also defined with spherical harmonics coefficients (SH) to represent view-dependent appearance and opacity for alpha value. The covariance matrix is valid only when it satisfies positive semi-definite, which is challenging to constrain during the optimization. Thus, the covariance matrix is decomposed into two learnable components, a quaternion $r$ for representing rotation and $s$ for representing scaling, to circumvent the positive semi-definite constraint similar to the configuration of an ellipsoid. $r$ and $s$ are transformed into rotation matrix and scaling matrix, respectively, and construct $\Sigma(r,s)$ as follows:
$\Sigma(r,s)=R(r)S(s)S(s)^{T}R(r)^{T},\vspace{0.2cm}$ (3)
where $R(r)$ is a rotation matrix given the rotation parameter $r$ and $S(s)$ is a scaling matrix from the scaling parameter $s$ [17]. These 3D Gaussians are projected to 2D space [48] to render 2D images with following 2D covariance matrix $\Sigma^{\prime}(r,s)$:
$\Sigma^{\prime}(r,s)=JW\Sigma(r,s)W^{T}J^{T},\vspace{0.2cm}$ (4)
where $J$ denotes the Jacobian of the affine approximation of the projective transformation, $W$ stands for the world-to-camera matrix. Each pixel value is computed by accumulating $N$ ordered projected 2D Gaussians overlaid on the each pixel with the formula:
$C=\sum_{i\in N}T_{i}c_{i}\alpha_{i}\hskip 5.0pt\text{ with }\hskip 5.0ptT_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),\vspace{0.2cm}$ (5)
$c_{i}$ is the color of each point, and $T_{i}$ is the transmittance. $\alpha_{i}\in[0,1]$ defined by $1-\exp^{-\sigma_{i}\delta_{i}}$ where $\sigma_{i}$ and $\delta_{i}$ are the density of the point and the interval along the ray respectively. For further details, please refer to the original paper [15].
3.2 Deblurring 3D Gaussians
3.2.1 Motivation
It is discussed in eq. 1 that the pixels in images get blurred due to defocusing, and this phenomenon is usually modeled through a convolution operation. According to the thin lens law [12], the scene points that lie at the focal distance of the camera make a sharp image at the imaging plane. On the other hand, any scene points that are not at the focal distance will make a blob instead of a sharp point on the imaging plane, and it produces a blurred image. If the separation from the focal distance of a scene point is large, it produces a blob of a large area, which corresponds to severe blur. This blurring effect is usually modeled by a 2D Gaussian function known as the point spread function (PSF). Correspondingly, an image captured by a camera is the result of the convolution of the actual image and the PSF. Through this convolution, which is the weighted summation of neighboring pixels, few pixels can affect the central pixel heavily depending on the weight. In other words, in the defocus imaging process, a pixel affects the intensity of neighboring pixels. This theoretical base provides us the motivation to build our deblurring 3D Gaussians framework. We assume that big sized 3D Gaussians cause the blur, while relatively smaller 3D Gaussians correspond to the sharp image. This is because those with greater dispersion are affected by more neighboring information as they are responsible for wider regions in image space, so they can represent the interference of the neighboring pixels. Whereas the fine details in the 3D scene can be better modeled through the smaller 3D Gaussians.
Figure 3: Comparison to normal Gaussian blur kernel. Top row: It shows the proposed method. $g$ is the Gaussian before the transformation, and $g^{\prime}$ is the Gaussian after the transformation. Since the different transformations can be applied to different Gaussian, ours can selectively blur the images depending on the scene; it can only blur the front parts of the scene. Bottom row: It shows a normal Gaussian blur kernel, which is not capable of handling different parts of the image differently, thereby uniformly blurring the entire image.
3.2.2 Modelling
Following the above mentioned motivation, we learn to deblur by transforming the geometry of the 3D Gaussians. The geometry of the 3D Gaussians is expressed through the covariance matrix, which can be decomposed into the rotation and scaling factors as mentioned in Eq. 3. Therefore, our target is to change the rotation and scaling factors of 3D Gaussians in such a way that we can model the blurring phenomenon. To do so, we have employed an MLP that takes the position $x$, rotation $r$, scale $s$, and viewing direction $v$ of 3D Gaussians as inputs, and outputs $(\delta r,\delta s)$, as given by:
$(\delta r,\delta s)=\mathcal{F_{\theta}}\Big{(}\gamma(x),r,s,\gamma(v)\Big{)},\vspace{0.2cm}$ (6)
where $\mathcal{F}_{\theta}$ denotes the MLP, and $\gamma$ denotes the positional encoding which is defined as:
$\gamma(p)=\big{(}\sin(2^{k}\pi p),\cos(2^{k}\pi p)\big{)}_{k=0}^{L-1},\vspace{0.2cm}$ (7)
where $L$ is the number of the frequencies, and the positional encoding is applied to each element of the vector $p$ [23].
The minima of these scaling factors $(\delta r,\delta s)$ are clipped to 1 and element-wisely multiplied to $r$ and $s$, respectively, to obtain the transformed attributes $r^{\prime}=r\cdot\delta r$ and $s^{\prime}=s\cdot\delta s$. With these transformed attributes, we can construct the transformed 3D Gaussians $G(x,r^{\prime},s^{\prime})$, which is optimized during training to model the scene blurriness. As $s^{\prime}$ is greater than or equal to $s$, each 3D Gaussian of $G(x,r^{\prime},s^{\prime})$ has greater statistical dispersion than the original 3D Gaussian $G(x,r,s)$. With the expanded dispersion of 3D Gaussian, it can represent the interference of the neighboring information which is a root cause of defocus blur. In addition, $G(x,r^{\prime},s^{\prime})$ can model the blurry scene more flexibly as per-Gaussian $\delta r$ and $\delta s$ are estimated. Defocus blur is spatially-varying, which implies different regions has different levels of blurriness. The scaling factors for 3D Gaussians that are responsible for a region with harsh defocus blur where various neighboring information in wide range is involved in, becomes bigger to better model high degree of blurriness. Those for 3D Gaussians on the sharp area, on the other hand, are closer to 1 so that they have smaller dispersion and do not represent the influence of the nearby information.
At the time of inference, we use $G(x,r,s)$ to render the sharp images. As mentioned earlier, we assume that multiplying two different scaling factors to transform the geometry of 3D Gaussians can work as blur kernel and convolution in eq. 1. Thus, $G(x,r,s)$ can produce the images with clean and fine details. It is worth noting that since any additional scaling factors are not used to render the images at testing time, $F_{\theta}$ is not activated so all steps required for the inference of Deblurring 3D-GS are identical to 3D-GS, which in turn enables real-time sharp image rendering. In terms of training, only forwarding $F_{\theta}$ and simple element-wise multiplication are extra cost.
3.2.3 Selective blurring
The proposed method can handle the training images arbitrarily blurred in various parts of the scene. Since we predict $(\delta r,\delta s)$ for each Gaussian, we can selectively enlarge the covariances of Gaussians where the parts in the training images are blurred. Such transformed per-Gaussian covariance is projected to 2D space and it can act as pixel-wise blur kernel at the image space. It is noteworthy that applying different shapes of blur kernels to different pixels plays a pivotal role in modeling scene blurriness since blurriness spatially varies. This flexibility enables us to effectively implement deblurring capability in 3D-GS. On the other hand, a naive approach to blurring the rendered image is simply to apply a Gaussian kernel. As shown in Fig. 3, this approach will blur the entire image, not blur pixel-wisely, resulting in blurring parts that should not be blurred for training the model. Even if learnable Gaussian kernel is applied, optimizing the mean and variance of the Gaussian kernel, a single type of blur kernel is limited in its expressivity to model the complicatedly blurred scene and is optimized to model the average blurriness of the scene from averaging loss function which fails to model blurriness morphing in each pixel. Not surprisingly, the Gaussian blur is a special case of the proposed method. If we predict one $(\delta s)$ for all 3D Gaussians, then it will similarly blur the whole image.
Algorithm 1 Add Extra Points
$\mathcal{P}$: Point cloud computed from SfM
$K$: Number of the neighboring points to find
$N_{p}$: Number of additional points to generate
$t_{d}$: Minimum required distance between new point and existing point
$P_{\text{add}}$GenerateRandomPoints($\mathcal{P},N_{p}$) $\triangleright$ Uniformly sample $N_{p}$ points
for each $p$ in $P_{\text{add}}$ do
     $\mathcal{P}_{\text{knn}}$FindNearestNeighbors($\mathcal{P},p,K$) $\triangleright$ Get $K$ nearest points of $p$ from $\mathcal{P}$
     $\mathcal{P}_{\text{valid}}\leftarrow$ CheckDistance($\mathcal{P}_{\text{knn}}$, $p$, $t_{d}$) $\triangleright$ Discard irrelevant neighbors
     if $|\mathcal{P}_{\text{valid}}|>0$  then
         $p_{c}\leftarrow$ LinearInterpolate($\mathcal{P}_{\text{valid}}$, $p$) $\triangleright$ Linearly interpolate neighboring colors
         AddToPointCloud($\mathcal{P},p,p_{c}$)
     end if
end for
3.3 Compensation for Sparse Point Cloud
3D-GS [15] constructs multiple 3D Gaussians from point clouds to model 3D scenes, and its reconstruction quality heavily relies on the initial point clouds. Point clouds are generally obtained from the structure-from-motion (SfM) [36], which extracts features from multi-view images and relates them to several 3D points. However, it can produce only sparse point clouds if the given images are blurry due to the challenge of feature extraction from blurry inputs. Even worse, if the scene has a large depth of field, which is prevalent in defocus blurry scenes, SfM hardly extracts any points that lie on the far end of the scene. To make a dense point cloud, we add additional points after $N_{st}$ iterations. $N_{p}$ points are sampled from an uniform distribution $U(\alpha,\beta)$ where $\alpha$ and $\beta$ are the minimum and maximum value of the position of the points from the existing point cloud, respectively. The color for each new point $p$ is assigned with the interpolated color $p_{c}$ from the nearest neighbors $\mathcal{P}_{\text{knn}}$ among the existing points using K-Nearest-Neigbhor (KNN) [29]. We discard the points whose distance to the nearest neighbor exceeds the distance threshold $t_{d}$ to prevent unnecessary points from being allocated to the empty space. The process of adding extra points to the given point cloud is summarized in algorithm 1. fig. 5 shows that a point cloud with additional points has a dense distribution of points to represent the objects.
Furthermore, 3D-GS [15] effectively manages the number of 3D Gaussians (the number of points) through periodic adaptive density control, densifying and pruning 3D Gaussians. To compensate for the sparsity of 3D Gaussians lying on the far end of the scene, we prune 3D Gaussians depending on their positions. As the benchmark Deblur-NeRF dataset [22] consists of only forward-facing scenes, the z-axis value of each point can be a relative depth from any viewpoints. As shown in fig. 4, we prune out less 3D Gaussians placed on the far edge of the scene to preserve more points located at the far plane, relying on the relative depth. Specifically, the pruning threshold $t_{p}$ is scaled by $w_{p}$ to 1.0 depending on the relative depth, and the lowest threshold is applied to the farthest point. Densifying more for far-plane-3D Gaussians is also an available option, but excessive compensation hinders modeling the scene blurriness, and it can further degrade real-time rendering as extra computational cost is required. We empirically found that flexible pruning was enough to compensate for sparse point cloud, considering the rendering speed. fig. 6 shows that relative depth-based pruning preserves the points lying on the far plane of the scene and leads to successful reconstruction of the objects on the far plane.
Figure 4: Comparison to pruning 3D Gaussians. Left: Given 3D Gaussians. Middle: Applying pruning method proposed by 3D-GS which removes 3D Gaussians with the single threshold ($t_{p}$). Right: Our pruning method which discards unnecessary 3D Gaussians with different thresholds based upon their depth.
4 Experiments
Table 1: Quantitative results on real defocus dataset. We color each cell as best and second best. *3D-GS achieved the highest FPS because it failed to model the blurry scene, abnormally small number of 3D Gaussians are in use which in turn raises FPS.
Cake Caps Cisco Coral Cupcake
[…]
NeRF [23] 24.42 0.7210 22.73 0.6312 20.72 0.7217 19.81 0.5658 21.88 0.6809
3D-GS [15] 20.16 0.5903 19.08 0.4355 20.01 0.6931 19.50 0.5519 21.53 0.6794
Deblur-NeRF [22] 26.27 0.7800 23.87 0.7128 20.83 0.7270 19.85 0.5999 22.26 0.7219
DP-NeRF [19] 26.16 0.7781 23.95 0.7122 20.73 0.7260 20.11 0.6107 22.80 0.7409
PDRF-10 [28] 27.06 0.8032 24.06 0.7102 20.68 0.7239 19.61 0.5894 22.95 0.7421
Ours 27.08 0.8076 24.68 0.7437 21.06 0.7409 20.24 0.5617 22.65 0.7477
Cups Daisy Sausage Seal Tools Average FPS
[…]
NeRF [23] 25.02 0.7581 22.74 0.6203 17.79 0.4830 22.79 0.6267 26.08 0.8523 22.40 0.6661 $<$ 1
3D-GS [15] 20.55 0.6459 20.96 0.6004 17.83 0.4718 22.25 0.5905 23.82 0.805 20.57 0.6064 $*{}^{*}$
Deblur-NeRF [22] 26.21 0.7987 23.52 0.6870 18.01 0.4998 26.04 0.7773 27.81 0.8949 23.46 0.7199 $<$ 1
DP-NeRF [19] 26.75 0.8136 23.79 0.6971 18.35 0.5443 25.95 0.7779 28.07 0.8980 23.67 0.7299 $<$ 1
PDRF-10 [28] 26.39 0.8066 24.49 0.7426 18.94 0.5686 26.36 0.7959 28.00 0.8995 23.85 0.7382 $<$ 1
Ours 26.26 0.8240 23.70 0.7268 18.74 0.5539 26.25 0.8197 27.58 0.9037 23.82 0.7430 620
Table 2: Quantitative results on synthetic defocus dataset. We color each cell as best and second best. *3D-GS achieved the highest FPS because it failed to model the blurry scene, abnormally small number of 3D Gaussians are in use which in turn raises FPS.
Cozyroom Factory Pool Tanabata Trolley Average FPS
[…]
NeRF [23] 30.03 0.8926 25.36 0.7847 27.77 0.7266 23.90 0.7811 22.67 0.7103 25.93 0.7791 $<$ 1
3D-GS [15] 30.09 0.9024 24.52 0.8057 20.14 0.4451 23.08 0.7981 22.26 0.7400 24.02 0.7383 789*
Deblur-NeRF [22] 31.85 0.9175 28.03 0.8628 30.52 0.8246 26.26 0.8517 25.18 0.8067 28.37 0.8527 $<$ 1
DP-NeRF [19] 32.11 0.9215 29.26 0.8793 31.44 0.8529 27.05 0.8635 26.79 0.8395 29.33 0.8713 $<$ 1
PDRF-10 [28] 32.29 0.9305 30.90 0.9138 30.97 0.8408 28.18 0.9006 28.07 0.8799 30.08 0.8931 $<$ 1
Ours 32.03 0.9269 29.89 0.9096 30.50 0.8344 27.56 0.9071 27.18 0.8755 29.43 0.8907 663
We compared our method against the state-of-the-art deblurrig approaches in neural rendering: Deblur-NeRF [22], DP-NeRF [20], PDRF [28] and original 3D Gaussians Splatting (3D-GS) [15]
4.1 Experimental Settings
We utilized PyTorch [27] to create our deblurring pipeline while maintaining the differentiable Gaussian rasterization 3D-GS [15]. We conducted 30k iterations of training. Adam optimizer [16] is used for optimization with the identical setting of 3D-GS. We set the learning rate for MLP at 1e-3, and the rest parameters are identical to 3D-GS. We use MLP with 3 hidden layers, each layer having 64 hidden units and ReLU activation. MLP is initialized with Xavier intialization [11]. For adding extra points to compensate the sparse point cloud, we set the addition start iteration $N_{st}$ to 2,500, the number of supplementing points $N_{p}$ to 100,000, the number of neighbors $K$ to 4, and the minimum distance threshold $t_{d}$ to 10. In terms of depth-based pruning, $t_{p}$ for pruning threshold is set to 5e-3 and 0.3 is used for $w_{p}$, the pruning threshold scaler. All the experiments were conducted on NVIDIA RTX 4090 GPU with 24GB memory.
We evaluated the performance on the benchmark Deblur-NeRF dataset [22] that includes both synthetic and real defocus blurred images. This dataset contains five scenarios in the synthetic defocus category and ten in the real defocus category. Synthetic defocus images are constructed using Blender [5]. By setting the aperture size and randomly selecting a focal plane between the closest and farthest depths, realistic defocus blur effects were produced. COLMAP [36, 37] is used to calculate the camera position of the blur and reference image in the real scene after choosing a big aperture to capture the defocus image.
Figure 5: Comparison on densifying point clouds during training. Left: Example training view. Middle: Point cloud at 5,000 training iterations without adding points. Bottom: Point cloud at 5,000 training iterations with adding extra points at 2,500 iterations.
4.2 Results and Comparisons
In this section, we provide the outcomes of our experiments, presenting a thorough analysis of both qualitative and quantitative results. Our evaluation framework encompasses a diverse set of metrics to show a comprehensive assessment of the experimental results. Primarily, we rely on established metrics such as the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Frames Per Second (FPS).
First of all, we show the transformation in the values of two Gaussian attributes (scale and rotation) to model the blur. The average values of these two attributes have been mentioned in Table 3. These values have been computed for all the scenes in real defocus dataset. It can be seen that during training, as compared to testing, 3D Gaussians of larger values of scale and rotation have been used to rasterize the scene. This indicates that larger values of scale and rotation are needed to adjust the 3D Gaussians to rasterize the blurred (training) images. While, in contrast, smaller values of these two attributes demonstrate that the smaller sized Gaussians are more suitable to model the fine details that are present in the sharp (testing) images.
Figure 6: Comparison on applying depth-based pruning. Top: Rendered image from the model with depth-dependent pruning. Bottom: Rendered image from the model with naive pruning as 3D-GS does.
As shown in  table 1 and  fig. 1 our method is on par with the state-of-the-art model in PSNR and achieve state-of-the-art performance evaluated under SSIM on the real defocus dataset. At the same time, it can still enjoy real-time rendering, with noticeable FPS, compared to other deblurring models. Also,  table 2 shows the quantitative result on the synthetic defocus dataset. FPS of 3D-GS is abnormally high because it fails to model the blurry scene, thus only small number of 3D Gaussians are created. fig. 7 and fig. 8 show the qualitative results on real and synthetic defocus dataset, respectively. We can see that ours can produce sharp and fine details, while 3D-GS fail to reconstruct those details.
4.3 Ablation Study
4.3.1 Extra points allocation
In this section, we evaluate the effect of adding extra points to the sparse point cloud. As shown in fig. 5, directly using sparse point cloud without any point densification only represents the objects with a small number of points or totally fails to model the tiny objects. Meanwhile, in case extra points are added to the point cloud, points successfully represent the objects densely. Also, the quantitative result is presented in table 4. It shows assigning valid color features to the additional points is important to deblur the scene and reconstruct the fine details.
Figure 7: Qualitative results on real defocus blur dataset.
Figure 8: Qualitative results on synthetic defocus blur dataset.
4.3.2 Depth-based pruning
We conduct an ablation study on depth-based pruning. To address excessive sparsity of point cloud at the far plane, we prune the points on the far plane less to maintain more numbers of points. table 5 shows there is severe rendering quality degradation if naive pruning, prunes the points with the same threshold from near plane to far plane, is used. In addition, fig. 6 shows a failure to reconstruct objects at the far plane when naive pruning is used, while objects lying on the near-end of the scene are well reconstructed. Meanwhile, ours, with depth-based pruning, can render clean objects on both near and far planes.
Table 3: Transformation of two attributes (scale and rotaion) of 3D Gaussians to model the blur.
train test
scale 0.87 0.85
rotation 0.26 0.25
Table 4: Ablation study on adding the extra points. w/ Random Colors stands for uniformly allocating points to the point cloud but color features are randomly initialized, rather than interpolated from neighboring points.
Methods PSNR $\uparrow$ SSIM $\uparrow$
w/o Extra Points 22.56 0.7016
w/ Random Colors 22.90 0.7089
w/ Extra Points 23.80 0.7430
Table 5: Ablation study on depth-depending pruning. Naive pruning stands for using naive points pruning from 3D-GS and Depth-based pruning stands for applying our depth-based pruning.
Methods PSNR $\uparrow$ SSIM $\uparrow$
Naive pruning 23.33 0.7296
Depth-based pruning 23.80 0.7430
5 Limitations & Future Works
NeRF-based deblurring methods [22, 20, 28], which are developed under the assumption of volumetric rendering, are not straightforwardly applicable to rasterization-based 3D-GS [15]. However, they can be compatible to rasterization by optimizing their MLP to deform kernels in the space of the rasterized image instead of letting MLP deform the rays and kernels in the world space. Although it is an interesting direction, it will incur additional costs for interpolating pixels and just implicitly transform the geometry of 3D Gaussians. Therefore, we believe that it will not be an optimal way to model scene blurriness using 3D-GS [15]. Another approach to deblur the defocus blur using 3D-GS [15] can be convolving the rasterized image with conventional grid blur kernels such as gaussian blur kernel [14]. With this approach however, the kernel expressivity is limited compared to the learnable kernels since the Gaussian blur kernel is a simple unimodal Gaussian distribution along each axis and is not learnable. Instead of a fixed Gaussian blur kernel, adopting a learnable grid blur kernel can also be another promising direction to address defocus blur. Nevertheless, as blurriness naturally varies spatially across the scene and each pixel, it is ideal to train blur kernels for every pixel from all training views. Thus, such method can incur high memory requirements, and its usage is limited if the rendering image is high resolution or lots of training images are provided.
6 Conclusion
We proposed Deblurring 3D-GS, the first defocus deblurring algorithm for 3D-GS. We adopted a small MLP that transforms the 3D Gaussians to model the scene blurriness. We also further facilitate deblurring by densifying sparse point clouds from blurry input images through additional point allocation, which uniformly distributes points in the scene and assigns color features using the K-Nearest-Neighbor algorithm. Also, as we applied depth-based pruning instead of naive pruning, which 3D-GS has adopted, we could preserve more points at the far edge of the scene where SfM usually struggles to extract features and fails to generate enough points. Through extensive experiments, we validated that our method can deblur the defocus blur while still enjoying the real-time rendering with FPS $>$ 200. This is because we use the MLP only during the training time, and the MLP is not involved in the inference stage, keeping the inference stage identical to the 3D-GS. Our method achieved state-of-the-art performance or performed comparably with current cutting-edge models evaluated under different metrics.

### v2

@@
 Qualitative results are available at https://benhenryl.github.io/Deblurring-3D-Gaussian-Splatting/
+Keywords:
+Neural Radiance Fields Deblurring Real-time rendering
 1 Introduction
@@
 Figure 1: Performance comparison to state-of-the-art deblurring NeRFs.
-Ours achieved a fast rendering speed (200 fps vs.
-1 fps) while maintaining competitive rendered image quality (x-axis is represented in log scale).
+Ours achieved a fast rendering speed ($>$ 800 FPS vs.
+1 FPS) while maintaining competitive rendered image quality (the x-axis is represented in log scale).
 Recently, 3D Gaussian Splatting (3D-GS) [14] has gained significant attention, demonstrating a capability to produce high-quality images at a remarkably fast rendering speed.
-It combines a large number of colored 3D Gaussians to represent 3D scenes with a differentiable splatting-based rasterization, substituting NeRF’s time-demanding volumetric rendering.
-The rendering process projects 3D Gaussian points onto the 2D image planes, and the positions, sizes, rotations, colors, and opacities of these Gaussians are adjusted via gradient-based optimization to better capture the 3D scenes.
-3D-GS leverages rasterization, which can be significantly more efficient than volume rendering techniques on modern graphics hardware, thereby enabling rapid real-time rendering.
+Substituting NeRF’s time-demanding volumetric rendering, it combines a large number of colored 3D Gaussians to represent 3D scenes with a differentiable splatting-based rasterization, which can be significantly more efficient than volume rendering techniques on modern graphics hardware, thereby enabling rapid real-time rendering.
 Expanding on the impressive capabilities of 3D-GS, we aim to further improve its robustness and versatility for more realistic settings, especially those involving blurring effects.
 Several approaches have attempted to handle the blurring issues in the recent NeRF literature [22, 20, 5, 39, 40].
-The pioneering work is Deblur-NeRF [22], which renders sharp images from the neural radiance fields and uses an extra multi-layer perceptron (MLP) to produce the blur kernels.
+The pioneering work is Deblur-NeRF [22], which renders sharp images from images with defocus blur or camera motion blur using an extra multi-layer perceptron (MLP) to produce the blur kernels.
 DP-NeRF [20] constrains neural radiance fields with two physical priors derived from the actual blurring process to reconstruct clean images.
@@
 All works mentioned above have been developed under the assumption of volumetric rendering, which is not straightforwardly applicable to rasterization-based 3D-GS.
-Another line of works [41, 6], though not dependent on volume rendering, only address a specific type of blur, i.e., camera motion blur, and are not valid for mitigating the defocus blur.
-In this work, we propose Deblurring 3D-GS, the first defocus deblurring algorithm for 3D-GS, which is well aligned with rasterization and thus enables real-time rendering.
+Another line of works [39, 5, 18] though not dependent on volume rendering, only address a single specific type of blur, i.e., either camera motion blur or defocus blur, and are not valid for mitigating the both types of blur.
+In this work, we propose Deblurring 3D-GS, the first deblurring algorithm for 3D-GS, which is well aligned with rasterization and thus enables real-time rendering.
 To do so, we modify the covariance matrices of 3D Gaussians to model the blurriness.
-Specifically, we employ a small MLP, which manipulates the covariance of each 3D Gaussian to model the scene blurriness.
+Specifically, we employ a small MLP, which manipulates the covariance mean of each 3D Gaussian to model the scene blurriness.
 As blurriness is a phenomenon that is based on the intermingling of the neighboring pixels, our Deblurring 3D-GS simulates such an intermixing during the training time.
 To this end, we designed a framework that utilizes an MLP to learn the variations in different attributes of 3D Gaussians.
-These small variations are multiplied to the original values of the attributes, which in turn determine the updated shape of the resulting Gaussians.
-During the inference time, we render the scene using only the original components of 3D-GS without any additional offsets from MLP; thereby, 3D-GS can render sharp images because each pixel is free from the intermingling of nearby pixels.
+These small variations are multiplied or added to the original values of the attributes, which in turn determine the updated shape of the resulting Gaussians.
+During the inference time, we render the scene using only the original components of 3D-GS without any additional outputs from the MLP; thereby, 3D-GS can render sharp images because each pixel is free from the intermingling of nearby pixels.
 Further, since the MLP is not activated during the inference time, it can still enjoy real-time rendering similar to the 3D-GS while it can reconstruct fine and sharp details from the blurry images.
 Figure 2: Our method’s overall workflow.
-$\gamma(\cdot)$ denotes positional encoding, $\odot$ denotes hadamard product, and $x$, $r$, $s$ stand for position, quaternion, and scaling of 3D Gaussian respectively.
-Given a set of 3D Gaussians $G$, we extract $x$, $r$, $s$ from each 3D Gaussian.
-Viewing direction $v$ from the dataset and $x$ are positionally encoded and then fed to $F_{\theta}$, a small MLP parameterized with $\theta$, which yields offsets $\delta r$ and $\delta s$ for each Gaussian.
-These $\delta r$ and $\delta s$ are element-wisely multiplied to their respective Gaussians’s $r$ and $s$ respectively, and these computed attributes are used to construct transformed 3D Gaussians $G^{\prime}(x,r\cdot\delta r,s\cdot\delta s)$.
-These transformed Gaussians $G^{\prime}$ are differentiably rasterized to render blurred image during training time.
-However, for inference stage, $G$ is directly fed to differentiable rasterizer, without involving any MLP, to render sharp image.
+$\gamma(\cdot)$ denotes positional encoding, $\odot$ and $\oplus$ denotes hadamard product and averaging operation for each, and $x$, $r$, $s$ stand for position, quaternion, and scaling of 3D Gaussian respectively.
+$\otimes$ is an operator that implements $\delta r\odot r$, $\delta s\odot s$, and $\delta x+x$.
+Dotted arrows and dashed arrows describe the pipeline for modeling camera motion blur and modeling defocus blur, respectively at training time.
+Solid arrows show the process of rendering sharp images at the inference time.
+More details are explained at  Sec.
+3.2.
 3D-GS [14] models a 3D scene from a sparse point cloud, which is usually obtained from the structure-from-motion (SfM) [34].
@@
 Due to this excessive sparsity in the point cloud constructed from set of blurry images, existing methods, including 3D-GS [14], that rely on point clouds fail to reconstruct the scene with fine details.
-To compensate for this excessive sparsity, we propose to add extra points with valid color features to the point cloud using N-nearest-neighbor interpolation [29].
+To compensate for this excessive sparsity, we propose to add extra points with valid color features to the point cloud using K-nearest-neighbor interpolation [29].
 In addition, we prune Gaussians based on their position to keep more Gaussians on the far plane.
 A variety of experiments have been conducted on the benchmark, and the results have revealed the effectiveness of our approach for deblurring.
-Tested under different evaluation matrices, our method achieves state-of-the-art rendering quality or performs on par with the currently leading models while achieving significantly faster rendering speed ($>$ 200 FPS)
+Tested under different evaluation matrices, our method achieves state-of-the-art rendering quality or performs on par with the currently leading models while achieving significantly faster rendering speed ($>$ 800 FPS)
 To sum up, our contributions are the following:
-  • We propose the first real-time rendering-enabled defocus deblurring framework using 3D-GS.
-  • We propose a novel technique that manipulates the covariance matrix of each 3D Gaussian differently to model spatially changing blur using a small MLP.
+  • We propose the first real-time rendering deblurring framework using 3D-GS.
+  • We propose a novel technique that manipulates the covariance matrix and mean of each 3D Gaussian differently to model spatially changing blur using a small MLP.
   • To compensate for sparse point clouds due to the blurry images, we propose a training technique that prunes and adds extra points with valid color features so that we can put more points on the far plane of the scene and harshly blurry regions.
-  • We achieve FPS $>$ 200 while accomplishing superior rendering quality or performing on par with the existing cutting-edge models under different metrics.
+  • We achieve FPS $>$ 800 while accomplishing superior rendering quality or performing on par with the existing cutting-edge models under different metrics.
 2 Related Works
 2.1 Image Deblurring
-It is common to observe that when we casually take pictures with optical imaging systems, some parts of scene appear blurred in the images.
+It is common to observe that when we casually take pictures with optical imaging systems, some parts of the scene appear blurred in the images.
 This blurriness is caused by a variety of factors, including object motion, camera shake, and lens defocusing [1, 33].
-When the image plane is separated from the ideal reference plane during the imaging process, out-of-focus or defocus blur happens.
-The degradation induced by defocus blur of an image is generally expressed as follows:
+The degradation induced by the blur of an image is generally expressed as follows:
 $g(x)=\sum_{s\in S_{h}}h(x,s)f(x)+n(x),~{}x\in S_{f},$ (1)
@@
 $S_{f}\subset\mathbb{R}^{2}$ is a support set of an image and $S_{h}\subset\mathbb{R}^{2}$ is a support set of a blur kernel or PSF [17].
-The first canonical approach to deconvolve an image is [33].
-It applies iterative minimization to an energy function to obtain a maximum likelihood approximation of the original image, using a known PSF (the blur kernel).
-Typically, the blurring kernel $K$ and the sharp image $I$ are unknown.
-The technique termed as image blind deblurring [42] involves recovering the latent sharp image given just one blurry image $B$.
-In this procedure, the deblurred image is obtained by first estimating a blurring kernel.
-Image blind deblurring is a long-standing and ill-posed problem in the field of image and vision.
-Numerous approaches to address the blurring issue have been suggested in the literature.
-Among them, traditional methods often construct deblurring as an optimization problem and rely on natural image priors [21, 26, 44, 47].
+Traditional methods often construct deblurring as an optimization problem and rely on natural image priors [21, 27, 41, 45].
 Conversely, the majority of deep learning-based techniques use convolutional neural networks (CNN) to map the blurry image with the latent sharp image directly [26, 44, 31].
-While series of studies have been actively conducted for image deblurring, they are mainly designed for deblurring 2D images and are not easily applicable to 3D scenes deblurring due to the lack of 3D view consistency.
+While a series of studies have been actively conducted for image deblurring, they are mainly designed for deblurring 2D images and are not easily applicable to 3D scenes deblurring due to the lack of 3D view consistency.
 2.2 Neural Radiance Fields
-Neural Radiance Fields (NeRF) are a potent method that has recently gained popularity for creating high-fidelity 3D scenes from 2D images.
-Realistic rendering from novel viewpoints is made possible by NeRF, which uses deep neural networks to encode volumetric scene features.
+Neural Radiance Fields (NeRF) is a potent method that has gained popularity for creating high-fidelity 3D scenes from 2D images, employing deep neural networks to encode volumetric scene features.
 To estimate density $\sigma\in[0,\infty)$ and color value $c\in[0,1]^{3}$ of a given point, a radiance field is a continuous function $f$ that maps a 3D location $x\in\mathbb{R}^{3}$ and a viewing direction $d\in\mathbb{S}^{2}$.
-This function has been parameterized by a multi-layer perceptron (MLP) [23], where the weights of MLP are optimized to reconstruct a series of input photos of a particular scene: $(c,\sigma)=f_{\theta}:\big{(}\gamma(x),\gamma(d)\big{)}$ Here, $\theta$ indicates the network weights, and $\gamma$ is the specified positional encoding applied to $x$ and $d$ [39].
+This function has been parameterized by a multi-layer perceptron (MLP) [23], where the weights of MLP are optimized to reconstruct a series of input photos of a particular scene: $(c,\sigma)=f_{\theta}:\big{(}\gamma(x),\gamma(d)\big{)}$.
+Here, $\theta$ indicates the network weights, and $\gamma$ is the specified positional encoding applied to $x$ and $d$ [37].
 To generate the images at novel views, volume rendering [6] is used, taking into account the volume density and color of points.
@@
 Numerous follow-up studies have been carried out to enhance NeRF’s rendering time to achieve real-time rendering.
-Many methods, such as grid-based approaches [40, 4, 38, 3, 8, 32], or those relying on hash [24, 2] adopt additional data structures to effectively reduce the size and number of layers of MLP and successfully improve the inference speed.
+Many methods, such as grid-based approaches [38, 4, 36, 3, 8, 32, 25], or those relying on hash [24, 2] adopt additional data structures to effectively reduce the size and number of layers of MLP and successfully improve the inference speed.
 However, they still fail to reach real-time view synthesis.
@@
 Deblur-NeRF [22] firstly suggests deblurring NeRF without any all-in-focus images during training.
-It employs an additional small MLP, which predicts per-pixel blur kernel.
-However, it has to render the neighboring pixels of the target pixel in advance to model the interference of the neighboring information, so it takes additional computation at the time of training.
+It employs an additional small MLP, which predicts per-pixel blur kernel to model defocus and camera motion blur.
 Though the inference stage does not involve the blur kernel estimation, it is no different from the training with regard to rendering time as it is based on volumetric rendering which takes several seconds to render a single image.
-DP-NeRF [19] improved Deblur-NeRF in terms of image quality, and PDRF [28] partially replaces MLP to the grid and ameliorates rendering speed, still they depend on volumetric rendering to produce images and are not free from the rendering cost.
-Other works [6, 41] are aimed at addressing another type of blur, camera motion blur, rather than solving the long rendering time.
+DP-NeRF [19] and PDRF [28] further improved Deblur-NeRF, still they depend on volumetric rendering and are not free from the rendering cost.
+Other approaches [5, 39, 18] are bounded to addressing only one type of blur, either camera motion blur or defocus blur, and not aimed at solving the long rendering time.
 While these deblurring NeRFs successfully produce clean images from the blurry input images, there is room for improvement in terms of rendering time.
@@
 Based on the 3D-GS [14], we generate 3D Gaussians, and each Gaussian is uniquely characterized by a set of the parameters, including 3D position $x$, opacity $\sigma$, and covariance matrix derived from quaternion $r$ scaling $s$.
-Each 3D Gaussian also contains spherical harmonics (SH) to represent view-dependent appearance.
+Every 3D Gaussian also contains spherical harmonics (SH) to represent view-dependent appearance.
 The input for the proposed method consists of camera poses and point clouds, which can be obtained through the structure from motion (SfM) [34], and a collection of images (possibly blurred).
-To deblur 3D Gaussians, we employ an MLP that takes the positions of the 3D Gaussians along with the $r$ and $s$ as inputs and outputs $\delta r$, and $\delta s$ which are the small scaling factors multiplied to $r$ and $s$, respectively.
-With new quaternion and scale, $r\cdot\delta r$ and $s\cdot\delta s$, the updated 3D Gaussians are subsequently fed to the tile-based rasterizer to produce the blurry images.
+We employ an MLP that takes $x_{j}$, $r_{j}$, and $s_{j}$ which are 3D position, quaternion, and scaling of $j$-th Gaussian, respectively as inputs to deblur a scene.
+In case of modeling defocus blur, the MLP yields $\delta r_{j}$ and $\delta s_{j}$ which are the small scaling factors multiplied to $r$ and $s$, respectively.
+With new quaternion and scale, $r_{j}\cdot\delta r_{j}$ and $s_{j}\cdot\delta s_{j}$, the updated 3D Gaussians are subsequently fed to the tile-based rasterizer to rasterize the defocus blurred images.
+To address camera motion blur, the MLP outputs $\{(\delta x_{j}^{(i)},\delta r_{j}^{(i)},\delta s_{j}^{(i)})\}_{i=1}^{M}$ where $M$ is the number of the auxiliary sets of 3D Gaussians representing the moments of camera movement and $\delta x_{j}^{(i)},\delta r_{j}^{(i)},\delta s_{j}^{(i)}$ are $i$-th predicted position offset, scaling factor for scale, and scaling factor for quaternion of $j$-th 3D Gaussian, respectively.
+Rasterizer produces $M$ images from $M$ different sets of 3D Gaussians and we average them to obtain camera motion blurred image.
 The overview of our method is shown in Fig.
@@
 It is discussed in Eq.
-1 that the pixels in images get blurred due to defocusing, and this phenomenon is usually modeled through a convolution operation.
-According to the thin lens law [12], the scene points that lie at the focal distance of the camera make a sharp image at the imaging plane.
-On the other hand, any scene points that are not at the focal distance will make a blob instead of a sharp point on the imaging plane, and it produces a blurred image.
-If the separation from the focal distance of a scene point is large, it produces a blob of a large area, which corresponds to severe blur.
-This blurring effect is usually modeled by a 2D Gaussian function known as the point spread function (PSF).
+1 that the pixels in images get blurred due to defocusing and camera motion, and this phenomenon is usually modeled through a convolution operation.
 Correspondingly, an image captured by a camera is the result of the convolution of the actual image and the PSF.
-Through this convolution, which is the weighted summation of neighboring pixels, few pixels can affect the central pixel heavily depending on the weight.
-In other words, in the defocus imaging process, a pixel affects the intensity of neighboring pixels.
-This theoretical base provides us the motivation to build our deblurring 3D Gaussians framework.
-We assume that big sized 3D Gaussians cause the blur, while relatively smaller 3D Gaussians correspond to the sharp image.
+Through convolution, which is the weighted summation of neighboring pixels, some pixels can affect the central pixel heavily depending on the weight.
+In other words, in the blurry imaging process, a pixel affects the intensity of neighboring pixels.
+This theoretical base motivates us to build our Deblurring 3D Gaussians framework.
+When handling defocus blur, we assume that big-sized 3D Gaussians cause the blur, while smaller 3D Gaussians correspond to the sharp image.
 This is because those with greater dispersion are affected by more neighboring information as they are responsible for wider regions in image space, so they can represent the interference of the neighboring pixels.
 Whereas the fine details in the 3D scene can be better modeled through the smaller 3D Gaussians.
-Figure 3: Comparison to normal Gaussian blur kernel.
-Top row: It shows the proposed method.
-$g$ is the Gaussian before the transformation, and $g^{\prime}$ is the Gaussian after the transformation.
-Since the different transformations can be applied to different Gaussian, ours can selectively blur the images depending on the scene; it can only blur the front parts of the scene.
-Bottom row: It shows a normal Gaussian blur kernel, which is not capable of handling different parts of the image differently, thereby uniformly blurring the entire image.
-3.2.2 Modelling
-Following the above mentioned motivation, we learn to deblur by transforming the geometry of the 3D Gaussians.
+In the case of camera motion blur, we implicitly model the camera movement during the camera exposure time.
+In detail, we generate multiple auxiliary sets of 3D Gaussians which represent the discrete moment of the movement, by shifting the positions of the existing set of 3D Gaussians, and simulate camera motion blur.
+More details are described in the supplementary material.
+3.2.2 Defocus blur modeling
+Following the aforementioned motivation, we learn to deblur by transforming the geometry of the 3D Gaussians.
 The geometry of the 3D Gaussians is expressed through the covariance matrix, which can be decomposed into the rotation and scaling factors as mentioned in Eq.
@@
 Therefore, our target is to change the rotation and scaling factors of 3D Gaussians in such a way that we can model the blurring phenomenon.
-To do so, we have employed an MLP that takes the position $x$, rotation $r$, scale $s$, and viewing direction $v$ of 3D Gaussians as inputs, and outputs $(\delta r,\delta s)$, as given by:
-$(\delta r,\delta s)=\mathcal{F_{\theta}}\Big{(}\gamma(x),r,s,\gamma(v)\Big{)},\vspace{0.2cm}$ (6)
+To do so, we have employed an MLP that takes the position $x_{j}$, rotation $r_{j}$, scale $s_{j}$ of $j$-th 3D Gaussian, and viewing direction $v$ as inputs, and outputs $(\delta r_{j},\delta s_{j})$, as given by:
+$(\delta r_{j},\delta s_{j})=\mathcal{F_{\theta}}\Big{(}\gamma(x_{j}),r_{j},s_{j},\gamma(v)\Big{)},\vspace{0.2cm}$ (6)
 where $\mathcal{F}_{\theta}$ denotes the MLP, and $\gamma$ denotes the positional encoding which is defined as:
@@
 where $L$ is the number of the frequencies, and the positional encoding is applied to each element of the vector $p$ [23].
-The minima of these scaling factors $(\delta r,\delta s)$ are clipped to 1 and element-wisely multiplied to $r$ and $s$, respectively, to obtain the transformed attributes $r^{\prime}=r\cdot\delta r$ and $s^{\prime}=s\cdot\delta s$.
-With these transformed attributes, we can construct the transformed 3D Gaussians $G(x,r^{\prime},s^{\prime})$, which is optimized during training to model the scene blurriness.
-As $s^{\prime}$ is greater than or equal to $s$, each 3D Gaussian of $G(x,r^{\prime},s^{\prime})$ has greater statistical dispersion than the original 3D Gaussian $G(x,r,s)$.
+Each scaling factor $(\delta r_{j},\delta s_{j})$ is scaled by $\lambda_{s}$, shifted by $(1-\lambda_{s})$ for optimization stability.
+Then the minima of them are clipped to 1.0 and element-wisely multiplied to $r_{j}$ and $s_{j}$, respectively, to obtain the transformed attributes as following:
+$\hat{r}_{j}=r_{j}\cdot\text{min}(1.0,~{}\lambda_{s}\delta r_{j}+(1-\lambda_{s})),$ (8)
+$\hat{s}_{j}=s_{j}\cdot\text{min}(1.0,~{}\lambda_{s}\delta s_{j}+(1-\lambda_{s})),$ (9)
+where function $\text{min}(\cdot,\cdot)$ returns smaller value.
+With these transformed attributes, we can construct the transformed 3D Gaussians $G(x_{j},\hat{r}_{j},\hat{s}_{j})$, which is optimized during training to model the scene blurriness.
+As $\hat{s}_{j}$ is greater than or equal to $s$, each 3D Gaussian of $G(x_{j},\hat{r}_{j},\hat{s}_{j})$ has greater statistical dispersion than the original 3D Gaussian $G(x_{j},r_{j},s_{j})$.
 With the expanded dispersion of 3D Gaussian, it can represent the interference of the neighboring information which is a root cause of defocus blur.
-In addition, $G(x,r^{\prime},s^{\prime})$ can model the blurry scene more flexibly as per-Gaussian $\delta r$ and $\delta s$ are estimated.
-Defocus blur is spatially-varying, which implies different regions has different levels of blurriness.
-The scaling factors for 3D Gaussians that are responsible for a region with harsh defocus blur where various neighboring information in wide range is involved in, becomes bigger to better model high degree of blurriness.
-Those for 3D Gaussians on the sharp area, on the other hand, are closer to 1 so that they have smaller dispersion and do not represent the influence of the nearby information.
-At the time of inference, we use $G(x,r,s)$ to render the sharp images.
+In addition, $G(x_{j},\hat{r}_{j},\hat{s}_{j})$ can model the blurry scene more flexibly as per-Gaussian $\delta r$ and $\delta s$ are estimated.
+Defocus blur is spatially varying, which implies different regions have different levels of blurriness.
+The scaling factors for 3D Gaussians that are responsible for a region with harsh defocus blur where various neighboring information in wide range is involved in, become bigger to better model a high degree of blurriness.
+Meanwhile, those for 3D Gaussians on the sharp area are closer to 1.0 so that they have smaller dispersion and do not represent the influence of the nearby information.
+Therefore, we can model defocus blur and rasterize defocus blurred image with $G(x_{j},\hat{r}_{j},\hat{s}_{j})$.
+At the time of inference, we use $G(x_{j},r_{j},s_{j})$ to render the sharp images.
 As mentioned earlier, we assume that multiplying two different scaling factors to transform the geometry of 3D Gaussians can work as blur kernel and convolution in Eq.
 1.
-Thus, $G(x,r,s)$ can produce the images with clean and fine details.
-It is worth noting that since any additional scaling factors are not used to render the images at testing time, $F_{\theta}$ is not activated so all steps required for the inference of Deblurring 3D-GS are identical to 3D-GS, which in turn enables real-time sharp image rendering.
-In terms of training, only forwarding $F_{\theta}$ and simple element-wise multiplication are extra cost.
-3.2.3 Selective blurring
-The proposed method can handle the training images arbitrarily blurred in various parts of the scene.
-Since we predict $(\delta r,\delta s)$ for each Gaussian, we can selectively enlarge the covariances of Gaussians where the parts in the training images are blurred.
-Such transformed per-Gaussian covariance is projected to 2D space and it can act as pixel-wise blur kernel at the image space.
-It is noteworthy that applying different shapes of blur kernels to different pixels plays a pivotal role in modeling scene blurriness since blurriness spatially varies.
-This flexibility enables us to effectively implement deblurring capability in 3D-GS.
-On the other hand, a naive approach to blurring the rendered image is simply to apply a Gaussian kernel.
-As shown in Fig.
-3, this approach will blur the entire image, not blur pixel-wisely, resulting in blurring parts that should not be blurred for training the model.
-Even if learnable Gaussian kernel is applied, optimizing the mean and variance of the Gaussian kernel, a single type of blur kernel is limited in its expressivity to model the complicatedly blurred scene and is optimized to model the average blurriness of the scene from averaging loss function which fails to model blurriness morphing in each pixel.
-Not surprisingly, the Gaussian blur is a special case of the proposed method.
-If we predict one $(\delta s)$ for all 3D Gaussians, then it will similarly blur the whole image.
+Thus, $G(x_{j},r_{j},s_{j})$ can produce the images with clean and fine details.
+It is worth noting that since any additional scaling factors are not used to render the images at testing time, $F_{\theta}$ is not activated, so all steps required for the inference of Deblurring 3D-GS are identical to 3D-GS, which in turn enables real-time sharp image rendering.
+3.2.3 Camera motion blur modeling
+We model camera motion blur with additional sets of 3D Gaussians.
+We adjust the geometry of each 3D Gaussian to simulate blur at training time, akin to defocus blur modeling.
+However, unlike defocus blur, camera motion blur occurs due to the physical movement of a camera.
+Every moment when the light hits the camera sensor, camera movement during the exposure time makes light intensities from multiple sources intermixed.
+We model such a phenomenon by adding small offsets to the position of each 3D Gaussian, and produce additional sets of 3D Gaussians to implicitly represent camera shake, and average clean images from different moments to simulate camera motion blur.
+Specifically, we first slightly change Eq.
+6 to compute additional output $\delta x$, the offset for the position $x_{j}$ of a Gaussian, from $F_{\theta}$ as:
+$\{(\delta x_{j}^{(i)},\delta r_{j}^{(i)},\delta s_{j}^{(i)})\}_{i=1}^{M}=\mathcal{F_{\theta}}\Big{(}\gamma(x_{j}),r_{j},s_{j},\gamma(v)\Big{)},\vspace{0.2cm}$ (10)
+where $M$ is the number of additional sets of 3D Gaussian to model the moments during the camera movement, $\delta x_{j}^{(i)}$ is $i$-th predicted position offset of $j$-th Gaussian.
+$\delta r_{j}^{(i)}~{}\text{and}~{}\delta s_{j}^{(i)}$ are scaled, shifted, and clipped in the same manner to Eq.
+8 and Eq.
+9 for each.
+Then we construct extra $M$ sets of 3D Gaussians $\{\{(\hat{x}_{j}^{(i)},\hat{r}_{j}^{(i)},\hat{s}_{j}^{(i)})\}_{i=1}^{M}\}_{j=1}^{N_{G}}$ by shifting the positions and changing the geometry of the existing set of 3D Gaussians, where $N_{G}$ is the number of the current 3D Gaussians, $\hat{x}_{j}^{(i)}$ stands for the shifted position with scaled $\delta x_{j}^{(i)}$ by $\lambda_{p}$.
+$\hat{x}_{j}^{(i)}=x_{j}+\lambda_{p}\delta x_{j}^{(i)}$, and $\hat{r}_{j}^{(i)}=r_{j}\cdot\delta r_{j}^{(i)}~{}\text{and}~{}\hat{s}_{j}^{(i)}=s_{j}\cdot\delta s_{j}^{(i)}$ are computed in a similar way to compute them for defocus blur deblurring.
+Each set corresponds to 3D Gaussians observed from different camera viewpoints and we rasterize $M$ clean images from $M$ different sets of 3D Gaussians and then average them to obtain a single camera motion blurred image $I_{b}$ at the training time as following:
+$I_{b}=\frac{1}{M}\sum_{i=1}^{M}I_{i},\quad I_{i}=\texttt{Rasterize}(\{G(\hat{x}_{j}^{(i)},\hat{r}_{j}^{(i)},\hat{s}_{j}^{(i)})\}_{j=1}^{N_{G}}),$ (11)
+where $I_{i}$ is a clean image generated by the predicted deltas.
+Deblurring camera motion blur also does not require any MLP forwarding and rendering multiple images at the inference time as $G(x_{j},r_{j},s_{j})$ learns the latent clean image.
+Thus, we can still enjoy rendering clean images from camera motion blurred input images in real-time manners, just like defocus blur deblurring.
 Algorithm 1 Add Extra Points
@@
 Point clouds are generally obtained from the structure-from-motion (SfM) [35], which extracts features from multi-view images and relates them to several 3D points.
-However, it can produce only sparse point clouds if the given images are blurry due to the challenge of feature extraction from blurry inputs.
+However, it can produce only sparse point clouds if the given images are blurry.
 Even worse, if the scene has a large depth of field, which is prevalent in defocus blurry scenes, SfM hardly extracts any points that lie on the far end of the scene.
-To make a dense point cloud, we add additional points after $N_{st}$ iterations.
-$N_{p}$ points are sampled from an uniform distribution $U(\alpha,\beta)$ where $\alpha$ and $\beta$ are the minimum and maximum value of the position of the points from the existing point cloud, respectively.
+To make a dense point cloud, we add extra points after $N_{st}$ iterations.
+$N_{p}$ points are sampled from a uniform distribution $U(\alpha,\beta)$ where $\alpha$ and $\beta$ are the minimum and maximum value of the position of the points from the existing point cloud, respectively.
 The color for each new point $p$ is assigned with the interpolated color $p_{c}$ from the nearest neighbors $\mathcal{P}_{\text{knn}}$ among the existing points using K-Nearest-Neigbhor (KNN) [29].
 We discard the points whose distance to the nearest neighbor exceeds the distance threshold $t_{d}$ to prevent unnecessary points from being allocated to the empty space.
-The process of adding extra points to the given point cloud is summarized in algorithm 1.
+The process of adding supplementary points to the given point cloud is summarized in Algorithm 1.
 Fig.
 3 shows that a point cloud with additional points has a dense distribution of points to represent the objects.
-Furthermore, 3D-GS [15] effectively manages the number of 3D Gaussians (the number of points) through periodic adaptive density control, densifying and pruning 3D Gaussians.
+Figure 3: Comparison on densifying point clouds during training.
+Left: Example training view.
+Middle: Point cloud at 5,000 training iterations without adding points.
+Right: Point cloud at 5,000 training iterations with adding extra points at 2,500 iterations.
+Furthermore, 3D-GS [14] effectively manages the number of 3D Gaussians through periodic adaptive density control, densifying and pruning 3D Gaussians.
 To compensate for the sparsity of 3D Gaussians lying on the far end of the scene, we prune 3D Gaussians depending on their positions.
-As the benchmark Deblur-NeRF dataset [22] consists of only forward-facing scenes, the z-axis value of each point can be a relative depth from any viewpoints.
+As the benchmark Deblur-NeRF dataset [22] consists of only forward-facing scenes, the z-axis value of each point can be a relative depth from any viewpoint.
 As shown in Fig.
 4, we prune out less 3D Gaussians placed on the far edge of the scene to preserve more points located at the far plane, relying on the relative depth.
-Specifically, the pruning threshold $t_{p}$ is scaled by $w_{p}$ to 1.0 depending on the relative depth, and the lowest threshold is applied to the farthest point.
-Densifying more for far-plane-3D Gaussians is also an available option, but excessive compensation hinders modeling the scene blurriness, and it can further degrade real-time rendering as extra computational cost is required.
-We empirically found that flexible pruning was enough to compensate for sparse point cloud, considering the rendering speed.
-fig.
-6 shows that relative depth-based pruning preserves the points lying on the far plane of the scene and leads to successful reconstruction of the objects on the far plane.
+Specifically, the pruning threshold $t_{p}$ is scaled by $\frac{1}{w_{p}}$ where $w_{p}$ is determined depending on the relative depth, and the lowest threshold is applied to the farthest point.
 Figure 4: Comparison to pruning 3D Gaussians.
 Left: Given 3D Gaussians.
-Middle: Applying pruning method proposed by 3D-GS which removes 3D Gaussians with the single threshold ($t_{p}$).
-Right: Our pruning method which discards unnecessary 3D Gaussians with different thresholds based upon their depth.
+Middle: Applying the pruning method proposed by 3D-GS which removes 3D Gaussians with the single threshold ($t_{p}$).
+Right: Our pruning method that discards unnecessary 3D Gaussians with different thresholds based on their depth.
 4 Experiments
-Table 1: Quantitative results on real defocus dataset.
-We color each cell as best and second best.
-*3D-GS achieved the highest FPS because it failed to model the blurry scene, abnormally small number of 3D Gaussians are in use which in turn raises FPS.
+We compared our method against the state-of-the-art deblurring approaches in neural rendering: Deblur-NeRF [22], Sharp-NeRF [18], DP-NeRF [20], PDRF [28] and original 3D Gaussians Splatting (3D-GS) [14] and image-based deblurring which deblurs training images first using Restormer [43] and then trains 3D-GS with them.
+We evaluated the performance on the benchmark Deblur-NeRF dataset [22] that includes both synthetic and real images captured with either camera motion blur or defocus blur.
+4.1 Experimental Settings
+We use Adam optimizer [15] and set the learning rate for MLP to $31e-$ 3, that for the position of 3D Gaussians to $31.6e-$ 3.
+3D Gaussian pruning threshold ($t_{p}$) and densification threshold are $35e-$ 3 and $42e-$ 4 respectively, for the real defocus blur dataset and $21e-$ 2 and $45e-$ 4 for the real camera motion blur dataset.
+The rest hyperparameters are identical to those of 3D-GS.
+We use an MLP with a depth of 4 layers.
+The first 3 layers are shared for all deltas and features from the shared layers are fed to each of 3 single layer (i.e., 1 layer head for each delta) that produces $\delta x,\delta r$, and $\delta s$ respectively.
+All layers have 64 hidden units, adopt ReLU activation for non-linearity, and are initialized with Xavier initialization [11].
+Both $\lambda_{p}$ and $\lambda_{s}$ are set to $21e-$ 2.
+For adding extra points to compensate for the sparse point cloud, we set the addition start iteration $N_{st}$ to 2,500, the number of supplementing points $N_{p}$ is proportional to the extent of the point cloud, at most 200,000, further explained in the supplementary material.
+The number of neighbors $K$ is 4, and the minimum distance threshold $t_{d}$ is 2.
+In terms of depth-based pruning, the pruning threshold multiplier $w_{p}$ is set to 3.
+We set $M$ for camera motion deblurring to 5 and the total iteration for training is 20,000.
+All the experiments were conducted on NVIDIA RTX 4090 GPU.
+4.2 Results and Comparisons
+In this section, we provide the outcomes of our experiments, presenting a thorough analysis of both qualitative and quantitative results.
+Our evaluation framework encompasses a diverse set of metrics to show a comprehensive assessment of the experimental results.
+Primarily, we rely on established metrics such as the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Frames Per Second (FPS).
+As shown in  Tab.
+1 and  Fig.
+1 our method is on par with the state-of-the-art model in PSNR and achieve state-of-the-art performance evaluated under SSIM on the real defocus blur dataset.
+Tab.
+2 further shows that the proposed method attains state-of-the-art performance on real camera motion blur dataset, under all metrics.
+At the same time, the proposed method can still enjoy real-time rendering, with a noticeable FPS, while other deblurring models cannot.
+Fig.
+5 shows the qualitative results on real camera motion blur dataset.
+We can see that ours can produce sharp and fine details, though 3D-GS fails to reconstruct those details.
+The qualitative and quantitative results on the rest datasets, more experiments and ablation studies are delivered in supplementary materials.
+Table 1: Quantitative results on real defocus blur dataset. We color each cell as best and second best.
 Cake Caps Cisco Coral Cupcake
@@
 3D-GS [14] 20.16 0.5903 19.08 0.4355 20.01 0.6931 19.50 0.5519 21.53 0.6794
+Restormer [43] + 3D-GS 21.48 0.6271 21.80 0.5956 19.95 0.6778 20.63 0.6125 23.18 0.7523
 Deblur-NeRF [22] 26.27 0.7800 23.87 0.7128 20.83 0.7270 19.85 0.5999 22.26 0.7219
+Sharp-NeRF [18] 26.23 0.7799 23.98 0.7098 20.88 0.7269 20.07 0.5999 22.75 0.7376
 DP-NeRF [19] 26.16 0.7781 23.95 0.7122 20.73 0.7260 20.11 0.6107 22.80 0.7409
 PDRF-10 [28] 27.06 0.8032 24.06 0.7102 20.68 0.7239 19.61 0.5894 22.95 0.7421
-Ours 27.08 0.8076 24.68 0.7437 21.06 0.7409 20.24 0.5617 22.65 0.7477
+Ours 26.88 0.8026 24.50 0.7428 20.83 0.7321 19.78 0.6080 22.11 0.7344
 Cups Daisy Sausage Seal Tools Average FPS
@@
 NeRF [23] 25.02 0.7581 22.74 0.6203 17.79 0.4830 22.79 0.6267 26.08 0.8523 22.40 0.6661 $<$ 1
-3D-GS [15] 20.55 0.6459 20.96 0.6004 17.83 0.4718 22.25 0.5905 23.82 0.805 20.57 0.6064 $*{}^{*}$
+3D-GS [14] 20.55 0.6459 20.96 0.6004 17.83 0.4718 22.25 0.5905 23.82 0.805 20.57 0.6064 788
+Restormer [43] + 3D-GS 21.89 0.7160 21.17 0.6196 17.99 0.5060 25.05 0.7364 24.63 0.8235 21.78 0.6669 775
 Deblur-NeRF [22] 26.21 0.7987 23.52 0.6870 18.01 0.4998 26.04 0.7773 27.81 0.8949 23.46 0.7199 $<$ 1
+Sharp-NeRF [18] 25.34 0.7783 23.66 0.7094 18.77 0.5482 25.82 0.7769 27.98 0.9032 23.55 0.7242 $<$ 1
 DP-NeRF [19] 26.75 0.8136 23.79 0.6971 18.35 0.5443 25.95 0.7779 28.07 0.8980 23.67 0.7299 $<$ 1
 PDRF-10 [28] 26.39 0.8066 24.49 0.7426 18.94 0.5686 26.36 0.7959 28.00 0.8995 23.85 0.7382 $<$ 1
-Ours 26.26 0.8240 23.70 0.7268 18.74 0.5539 26.25 0.8197 27.58 0.9037 23.82 0.7430 620
-Table 2: Quantitative results on synthetic defocus dataset.
-We color each cell as best and second best.
-*3D-GS achieved the highest FPS because it failed to model the blurry scene, abnormally small number of 3D Gaussians are in use which in turn raises FPS.
+Ours 26.28 0.8235 23.54 0.7310 18.99 0.5705 26.18 0.8166 27.96 0.9098 23.71 0.7471 804
+Table 2: Quantitative results on real camera motion blur dataset tested under PSNR, SSIM and FPS. We color each cell as best and second best.
+Ball Basket Buick Coffee Decoration
+[…]
+NeRF [23] 24.08 0.6237 23.72 0.7086 21.59 0.6325 26.47 0.8064 22.39 0.6609
+3D-GS [14] 22.99 0.6206 23.11 0.6833 21.22 0.6519 23.53 0.6995 20.45 0.6239
+Restormer [43] + 3D-GS 23.85 0.6498 23.75 0.7208 21.42 0.6949 23.94 0.7235 20.98 0.6840
+Deblur-NeRF [22] 27.36 0.7656 27.67 0.8449 24.77 0.7700 30.93 0.8981 24.19 0.7707
+DP-NeRF [19] 27.20 0.7652 27.74 0.8455 25.70 0.7922 31.19 0.9049 24.31 0.7811
+PDRF-10 [28] 27.96 0.7365 28.82 0.8465 25.52 0.7742 31.55 0.8627 23.26 0.7164
+Ours 28.27 0.8233 28.42 0.8713 25.95 0.8367 32.84 0.9312 25.87 0.8540
+Girl Heron Parterre Puppet Stair Average FPS
+[…]
+NeRF [23] 20.07 0.7075 20.50 0.5217 23.14 0.6201 22.09 0.6093 22.87 0.4561 22.69 0.6347 $<$ 1
+3D-GS [14] 19.72 0.7031 19.26 0.4767 22.22 0.5813 22.18 0.6362 21.88 0.4789 21.66 0.6154 734
+Restormer [43] + 3D-GS 19.71 0.7151 19.68 0.5615 22.60 0.6364 22.19 0.6608 22.66 0.5735 22.08 0.6620 708
+Deblur-NeRF [22] 22.27 0.7976 22.63 0.6874 25.82 0.7597 25.24 0.7510 25.39 0.6296 25.63 0.7675 $<$ 1
+DP-NeRF [19] 23.33 0.8139 22.88 0.6930 25.86 0.7665 25.25 0.7536 25.59 0.6349 25.91 0.7751 $<$ 1
+PDRF-10 [28] 23.78 0.8120 22.90 0.6590 25.19 0.7233 25.06 0.7326 25.73 0.5722 25.98 0.7245 $<$ 1
+Ours 23.26 0.8390 23.14 0.7438 26.17 0.8144 25.67 0.8051 26.46 0.7050 26.61 0.8224 961
+Figure 5: Qualitative results on real camera motion blur dataset.
+5 Limitations & Future Works
+NeRF-based deblurring methods [22, 20, 28], which are developed under the assumption of volumetric rendering, are not easily applicable to rasterization-based 3D-GS [14].
+However, they can be compatible to rasterization by optimizing their MLP to deform kernels in the space of the rasterized image instead of letting MLP deform the rays and kernels in the world space.
+Although it is an interesting direction, it will incur additional costs for interpolating pixels and just implicitly transform the geometry of 3D Gaussians.
+Therefore, we believe that it will not be an optimal way to model scene blurriness using 3D-GS [14].
+6 Conclusion
+We proposed Deblurring 3D-GS, the first deblurring algorithm for 3D-GS.
+We adopted a small MLP that transforms the 3D Gaussians to model the scene blurriness.
+We also further facilitated deblurring by complementing more points on sparse point clouds.
+We validated that our method can deblur the scene while still enjoying the real-time rendering with FPS $>$ 800.
+This is because we use the MLP only during the training time, and the MLP is not involved in the inference stage, keeping the inference stage identical to the 3D-GS.
+Supplementary Materials
+Appendix A Additional Method Details
+A.0.1 Defocus blur modeling
+According to the thin lens law [12], the scene points that lie at the focal distance of the camera make a sharp image at the imaging plane.
+On the other hand, any scene points that are not at the focal distance will make a blob instead of a sharp point on the imaging plane, and it produces a defocus blurred image.
+If the separation from the focal distance of a scene point is large, it produces a blob of a large area, which corresponds to severe defocus blur.
+We assume that 3D Gaussians with greater dispersion (greater scale $s$) can represent the scene points not being located at the focal distance, while those with smaller scale $s$ can model the points placed at the focal distance.
+Table 3 shows that during training, as compared to testing, 3D Gaussians of larger values of scale have been used to rasterize the scene.
+This indicates that larger values of scale are needed to adjust the 3D Gaussians to rasterize the blurred (training) images.
+While, in contrast, smaller values of the attributes demonstrate that the smaller-sized Gaussians are more suitable to model the fine details that are present in the sharp (testing) images.
+Table 3: Scale transformation of 3D Gaussians to model the defocus blur measured on real defocus blur dataset.
+Train ($\hat{s}_{j}$) Test ($s_{j}$)
+Average 0.3232 0.3079
+A.0.2 Camera motion blur modeling
+Camera motion blur occurs due to the camera shake during the exposure time of a photograph.
+When a camera moves while capturing an image, the rays from different points within the scene strike different areas of the camera sensor at different moments and are intermixed, so the blurry final image is obtained [7].
+Consequently, the blurry final image is acquired as each point in the subject is captured slightly different region on the sensor.
+Meanwhile, images at all the instants when the rays hit the camera sensor are sharp, without any inherent blur.
+Therefore, we train the 3D Gaussians to model clean representations at various moments the lights reach the sensor, rasterize $M$ clean images from $M$ different moments, and then average them to simulate the camera movement during training time.
+Fig.
+6 shows the clean images rasterized at different moments and averaged blurry image when $M$ is 6.
+Figure 6: Camera motion blur modeling during training.
+Clean images at different moments of the camera movement during exposure time are rasterized and averaged to a single image to model the camera motion blur.
+A.0.3 Selective defocus blurring
+The proposed method can handle the training images arbitrarily blurred in various parts of the scene.
+Since we predict $(\delta s_{j})$ for each Gaussian, we can selectively enlarge the covariances of Gaussians where the parts in the training images are blurred.
+Such transformed per-Gaussian covariance is projected to 2D space and it can act as pixel-wise blur kernel at the image space.
+It is noteworthy that applying different shapes of blur kernels to different pixels plays a pivotal role in modeling scene blurriness since blurriness spatially varies.
+This flexibility enables us to effectively implement deblurring capability in 3D-GS [14].
+On the other hand, a naive approach to blurring the rendered image is simply to apply a Gaussian kernel.
+As shown in Fig.
+7, this approach will blur the entire image, not blur pixel-wisely, resulting in blurring parts that should not be blurred for training the model.
+Even if a learnable Gaussian kernel is applied, optimizing the mean and variance of the Gaussian kernel, a single type of blur kernel is limited in its expressivity to model the complicatedly blurred scene and is optimized to model the average blurriness of the scene from averaging loss function which fails to model blurriness morphing in each pixel.
+Not surprisingly, the Gaussian blur is a special case of the proposed method.
+If we predict one $(\delta s_{j})$ for all 3D Gaussians, then it will similarly blur the whole image.
+Fig.
+8 shows that the proposed method successfully deblur the defocus blur, while normal Gaussian blur kernel approaches fail.
+Moreover, transforming per-Gaussian covariance allows to adjust scene blurriness arbitrarily as shown in Fig.
+9.
+Figure 7: Comparison to normal Gaussian blur kernel.
+Top row: It shows the proposed method.
+$g$ is the Gaussian before the transformation, and $g^{\prime}$ is the Gaussian after the transformation.
+Since the different transformations can be applied to different Gaussian, ours can selectively blur the images depending on the scene; it can only blur the front parts of the scene.
+Bottom row: It describes a normal Gaussian blur kernel where $h$ is the Gaussian, and $h^{\prime}$ is the Gaussian after applying a normal Gaussian blur kernel.
+Simply applying a normal Gaussian blur kernel is not capable of handling different parts of the image differently, thereby uniformly blurring the entire image.
+Figure 8: Defocus-deblurred images with different sizes of normal Gaussian blur kernels and the proposed method.
+(A), (B), (C): 15$\times$15, 9$\times$9, and 5$\times$5 Gaussian blur kernel are in use to deblur, respectively, and the bottom row shows the visualization of each kernel whose values are inverted for better visibility (D): Proposed method which transforms geometry of each 3D Gaussian.
+Figure 9: Selective Gaussian blur adjustment.
+As delineated in  Fig.
+7, our methodology adeptly harnesses the $\delta r_{j},\delta s_{j}$ both emanating from compact Multi-Layer Perceptrons (MLP), enabling the inversion of Gaussian blur regions or the comprehensive modulation of overall blurriness and sharpness.
+With the Transformation of $\delta r_{j},\delta s_{j}$, our framework facilitates the nuanced blurring of proximal regions akin to A, as well as the deft blurring of distant locales akin to B.
+Furthermore, it offers the capability to manipulate the global blurriness or sharpness, exemplified by adjustments akin to C and D.
+A.0.4 Visualization
+Fig.
+10 visualizes the original and transformed 3D Gaussians for defocus blur.
+With a given view whose near plane is defocused, the transformed 3D Gaussians show larger scales than those of the original 3D Gaussians to model defocus blur on the near plane (blur-bordered images).
+Meanwhile, the transformed 3D Gaussians keep very similar shapes to the original ones for sharp objects in the far plane (red-bordered images).
+Fig.
+11 depicts point clouds of the original 3D Gaussians and transformed 3D Gaussians.
+The point cloud of the transformed 3D Gaussians exhibits the camera movements when it moves left to right.
+Figure 10: Gaussians visualization for defocus blur.
+Figure 11: Point cloud visualization for camera motion blur.
+A.0.5 Adding extra points
+The number of supplementing points $N_{p}$ is determined based on the extent of the point cloud as follows:
+$N_{p}=\text{min}(\frac{\text{prod}(Q_{max}-Q_{min})}{c^{3}},N_{max}),$ (12)
+where $\text{prod}(\cdot)$ returns a product of all values in a vector, $Q_{max}\in\mathbb{R}^{3}$ and $Q_{min}\in\mathbb{R}^{3}$ are the maximum and minimum values of the positions of the points in the point cloud along x-, y-, and z-axis, respectively, and $c$ is a constant for stability which is set to 1.1.
+$N_{max}$ is 200,000 which restricts the number of points to be added.
+Appendix B Additional Experiments
+B.0.1 More results
+We further evaluate the proposed method under Learned Perceptual Image Patch Similarity (LPIPS) metric as shown in Tab.
+4 and Tab.
+5 for real defocus blur and real camera motion blur dataset.
+Our method achieves state-of-the-art performance under LPIPS on both datasets.
+Also, we conducted experiments on the synthetic defocus blur (Tab.
+6, and Tab.
+8) and camera motion blur (Tab.
+7, and Tab.
+8) dataset.
+For the synthetic defocus blur dataset, we set the learning rate for the position of 3D Gaussians to $44.8e-$ 4.
+The learning rate for scale and rotation attribute of 3D Gaussians is set to 0.015, and 0.005 respectively, $\lambda_{s}$ to 0.005, and $\lambda_{p}$ to 0.001 for the synthetic camera motion blur dataset.
+The rest hyperparameters are identical to those for real blur datasets.
+Qualitative results of each dataset are illustrated in Fig.
+14 to Fig.
+16.
+Table 4: Quantitative results on real defocus blur dataset tested under LPIPS metric. We color each cell as best and second best.
+Cake Caps Cisco Coral Cupcake Cups Daisy Sausage Seal Tools Average
+NeRF [23] 0.2250 0.2801 0.1256 0.2155 0.2689 0.2315 0.2621 0.2789 0.2680 0.1547 0.2310
+3D-GS [14] 0.2082 0.4329 0.1781 0.3111 0.2081 0.3211 0.2629 0.2855 0.3057 0.1953 0.2920
+Deblur-NeRF [22] 0.1282 0.1612 0.0868 0.1160 0.1214 0.1271 0.1208 0.1796 0.1048 0.0610 0.1207
+DP-NeRF [19] 0.1267 0.1430 0.0840 0.0960 0.1178 0.1035 0.1075 0.1473 0.1026 0.0539 0.1082
+PDRF-10 [28] 0.1622 0.2854 0.0943 0.2335 0.1862 0.1370 0.1024 0.2126 0.1927 0.1395 0.1746
+Ours 0.1158 0.1491 0.0794 0.1310 0.0985 0.1028 0.0949 0.1405 0.0980 0.0577 0.1068
+Table 5: Quantitative results on real camera motion blur dataset tested under LPIPS metric. We color each cell as best and second best.
+Ball Basket Buick Coffee Decoration Girl Heron Parterre Puppet Stair Average
+NeRF [23] 0.3992 0.3223 0.3502 0.2896 0.3633 0.3196 0.4129 0.4046 0.3389 0.4868 0.3687
+3D-GS [14] 0.3505 0.2925 0.2839 0.3113 0.3141 0.2843 0.3328 0.3197 0.2607 0.3888 0.3240
+Deblur-NeRF [22] 0.2230 0.1481 0.1752 0.1244 0.1862 0.1687 0.2099 0.2161 0.1577 0.2102 0.1820
+DP-NeRF [19] 0.2088 0.1294 0.1405 0.1002 0.1639 0.1498 0.1914 0.1900 0.1505 0.1772 0.1602
+PDRF-10 [28] 0.2487 0.1241 0.1751 0.1424 0.2379 0.1828 0.2367 0.2639 0.1569 0.2430 0.2012
+Ours 0.1413 0.1155 0.0954 0.0676 0.0933 0.1011 0.1543 0.1206 0.0941 0.1123 0.1096
+Table 6: Quantitative results on synthetic defocus blur dataset under PSNR and SSIM metrics. We color each cell as best and second best.
 Cozyroom Factory Pool Tanabata Trolley Average FPS
@@
 NeRF [23] 30.03 0.8926 25.36 0.7847 27.77 0.7266 23.90 0.7811 22.67 0.7103 25.93 0.7791 $<$ 1
-3D-GS [15] 30.09 0.9024 24.52 0.8057 20.14 0.4451 23.08 0.7981 22.26 0.7400 24.02 0.7383 789*
+3D-GS [14] 30.09 0.9024 24.52 0.8057 20.14 0.4451 23.08 0.7981 22.26 0.7400 24.02 0.7383 789
 Deblur-NeRF [22] 31.85 0.9175 28.03 0.8628 30.52 0.8246 26.26 0.8517 25.18 0.8067 28.37 0.8527 $<$ 1
+Sharp-NeRF [18] 31.32 0.9133 28.67 0.8979 30.51 0.8264 24.95 0.8536 26.03 0.8498 28.30 0.8682 $<$ 1
 DP-NeRF [19] 32.11 0.9215 29.26 0.8793 31.44 0.8529 27.05 0.8635 26.79 0.8395 29.33 0.8713 $<$ 1
 PDRF-10 [28] 32.29 0.9305 30.90 0.9138 30.97 0.8408 28.18 0.9006 28.07 0.8799 30.08 0.8931 $<$ 1
-Ours 32.03 0.9269 29.89 0.9096 30.50 0.8344 27.56 0.9071 27.18 0.8755 29.43 0.8907 663
-We compared our method against the state-of-the-art deblurrig approaches in neural rendering: Deblur-NeRF [22], DP-NeRF [20], PDRF [28] and original 3D Gaussians Splatting (3D-GS) [15]
-4.1 Experimental Settings
-We utilized PyTorch [27] to create our deblurring pipeline while maintaining the differentiable Gaussian rasterization 3D-GS [15].
-We conducted 30k iterations of training.
-Adam optimizer [16] is used for optimization with the identical setting of 3D-GS.
-We set the learning rate for MLP at 1e-3, and the rest parameters are identical to 3D-GS.
-We use MLP with 3 hidden layers, each layer having 64 hidden units and ReLU activation.
-MLP is initialized with Xavier intialization [11].
-For adding extra points to compensate the sparse point cloud, we set the addition start iteration $N_{st}$ to 2,500, the number of supplementing points $N_{p}$ to 100,000, the number of neighbors $K$ to 4, and the minimum distance threshold $t_{d}$ to 10.
-In terms of depth-based pruning, $t_{p}$ for pruning threshold is set to 5e-3 and 0.3 is used for $w_{p}$, the pruning threshold scaler.
-All the experiments were conducted on NVIDIA RTX 4090 GPU with 24GB memory.
-We evaluated the performance on the benchmark Deblur-NeRF dataset [22] that includes both synthetic and real defocus blurred images.
-This dataset contains five scenarios in the synthetic defocus category and ten in the real defocus category.
-Synthetic defocus images are constructed using Blender [5].
-By setting the aperture size and randomly selecting a focal plane between the closest and farthest depths, realistic defocus blur effects were produced.
-COLMAP [36, 37] is used to calculate the camera position of the blur and reference image in the real scene after choosing a big aperture to capture the defocus image.
-Figure 5: Comparison on densifying point clouds during training.
-Left: Example training view.
-Middle: Point cloud at 5,000 training iterations without adding points.
-Bottom: Point cloud at 5,000 training iterations with adding extra points at 2,500 iterations.
-4.2 Results and Comparisons
-In this section, we provide the outcomes of our experiments, presenting a thorough analysis of both qualitative and quantitative results.
-Our evaluation framework encompasses a diverse set of metrics to show a comprehensive assessment of the experimental results.
-Primarily, we rely on established metrics such as the Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Frames Per Second (FPS).
-First of all, we show the transformation in the values of two Gaussian attributes (scale and rotation) to model the blur.
-The average values of these two attributes have been mentioned in Table 3.
-These values have been computed for all the scenes in real defocus dataset.
-It can be seen that during training, as compared to testing, 3D Gaussians of larger values of scale and rotation have been used to rasterize the scene.
-This indicates that larger values of scale and rotation are needed to adjust the 3D Gaussians to rasterize the blurred (training) images.
-While, in contrast, smaller values of these two attributes demonstrate that the smaller sized Gaussians are more suitable to model the fine details that are present in the sharp (testing) images.
-Figure 6: Comparison on applying depth-based pruning. Top: Rendered image from the model with depth-dependent pruning. Bottom: Rendered image from the model with naive pruning as 3D-GS does.
-As shown in  table 1 and  fig.
-1 our method is on par with the state-of-the-art model in PSNR and achieve state-of-the-art performance evaluated under SSIM on the real defocus dataset.
-At the same time, it can still enjoy real-time rendering, with noticeable FPS, compared to other deblurring models.
-Also,  table 2 shows the quantitative result on the synthetic defocus dataset.
-FPS of 3D-GS is abnormally high because it fails to model the blurry scene, thus only small number of 3D Gaussians are created.
-fig.
-7 and fig.
-8 show the qualitative results on real and synthetic defocus dataset, respectively.
-We can see that ours can produce sharp and fine details, while 3D-GS fail to reconstruct those details.
-4.3 Ablation Study
-4.3.1 Extra points allocation
-In this section, we evaluate the effect of adding extra points to the sparse point cloud.
-As shown in fig.
-5, directly using sparse point cloud without any point densification only represents the objects with a small number of points or totally fails to model the tiny objects.
-Meanwhile, in case extra points are added to the point cloud, points successfully represent the objects densely.
-Also, the quantitative result is presented in table 4.
-It shows assigning valid color features to the additional points is important to deblur the scene and reconstruct the fine details.
-Figure 7: Qualitative results on real defocus blur dataset.
-Figure 8: Qualitative results on synthetic defocus blur dataset.
-4.3.2 Depth-based pruning
+Ours 31.97 0.9275 29.16 0.9089 31.31 0.8580 27.54 0.9083 27.55 0.8858 29.51 0.8977 798
+Table 7: Quantitative results on synthetic camera motion blur dataset under PSNR and SSIM metrics. We color each cell as best and second best.
+Cozyroom Factory Pool Tanabata Trolley Average FPS
+[…]
+NeRF [23] 25.66 0.7941 19.32 0.4563 30.45 0.8354 22.22 0.6807 21.25 0.6370 23.78 0.6807 $<$ 1
+3D-GS [14] 25.59 0.8076 18.11 0.4179 25.63 0.6326 21.35 0.6686 20.56 0.6257 22.25 0.6305 1019
+Deblur-NeRF [22] 32.08 0.9261 25.60 0.7750 31.61 0.8682 27.11 0.8640 27.45 0.8632 28.77 0.8593 $<$ 1
+DP-NeRF [19] 32.65 0.9317 25.91 0.7787 31.96 0.8768 27.61 0.8748 28.03 0.8752 29.23 0.8674 $<$ 1
+PDRF-10 [28] 31.90 0.9321 26.56 0.8102 31.29 0.8657 28.21 0.8952 28.48 0.8956 29.29 0.8798 $<$ 1
+Ours 31.45 0.9222 24.01 0.7333 31.87 0.8829 27.01 0.8807 26.88 0.8710 28.24 0.8580 919
+Table 8: Quantiative results on synthetic camera motion blur dataset tested under LPIPS metric. We color each cell as best and second best.
+Defocus blur Camera motion blur
+Cozy2room Factory Pool Tanabata Wine Average Cozy2room Factory Pool Tanabata Wine Average
+NeRF [23] 0.0885 0.2351 0.3340 0.2142 0.2799 0.2303 0.2288 0.5304 0.1932 0.3653 0.3633 0.3362
+3D-GS  [14] 0.0692 0.1842 0.5094 0.1710 0.2281 0.2323 0.1645 0.4958 0.2632 0.3235 0.3390 0.3172
+Deblur-NeRF [22] 0.0481 0.1127 0.1901 0.0995 0.1436 0.1188 0.0477 0.2687 0.1246 0.1228 0.1363 0.1400
+DP-NeRF [19] 0.0386 0.1035 0.1563 0.0779 0.1170 0.0987 0.0355 0.2494 0.0908 0.1033 0.1129 0.1184
+PDRF-10 [28] 0.0518 0.1066 0.1893 0.0819 0.1210 0.1101 0.0448 0.1499 0.1345 0.1025 0.0939 0.1051
+Ours 0.0424 0.1079 0.1587 0.0821 0.1163 0.1015 0.0367 0.2326 0.0751 0.0785 0.1028 0.1051
+B.0.2 Depth-based pruning
 We conduct an ablation study on depth-based pruning.
 To address excessive sparsity of point cloud at the far plane, we prune the points on the far plane less to maintain more numbers of points.
-table 5 shows there is severe rendering quality degradation if naive pruning, prunes the points with the same threshold from near plane to far plane, is used.
+Tab.
+9 shows our depth-based pruning can preserve more points located at the far plane which leads to better reconstruction quality than naive pruning, which prunes the points with a single threshold regardless of the positions of the points.
 In addition, Fig.
@@
 Meanwhile, ours, with depth-based pruning, can render clean objects on both near and far planes.
-Table 3: Transformation of two attributes (scale and rotaion) of 3D Gaussians to model the blur.
-train test
-scale 0.87 0.85
-rotation 0.26 0.25
+Table 9: Ablation study on depth-depending pruning. Naive pruning stands for using naive points pruning from 3D-GS and Depth-based pruning stands for applying our depth-based pruning.
+Methods PSNR $\uparrow$ SSIM $\uparrow$
+Naive pruning 23.52 0.7394
+Depth-based pruning 23.73 0.7474
+Figure 12: Comparison on applying depth-based pruning. Top row: Rendered image from the model with depth-dependent pruning. Bottom row: Rendered image from the model with naive pruning as 3D-GS does.
+B.0.3 Ablation study on $M$
+We run an ablation study on the hyperparameter $M$, the number of the moments to be averaged.
+As shown in Tab.
+10 our method shows similar performance when $5\leq M\leq 10$.
+Considering the increasing training time with higher $M$, we set $M$ to 5.
+Table 10: Ablation on $M$ under real camera motion blur dataset.
+PSNR Training time PSNR Training time
+$M=2$ 24.81 $\approx$ 8.0 mins $M=10$ 26.62 $\approx$ 22.4 mins
+$M=5$ 26.66 $\approx$ 13.4 mins $M=20$ 25.81 $\approx$ 43.4 mins
+B.0.4 Ablation study on extra points allocation
+In this section, we evaluate the effect of adding extra points to the sparse point cloud.
+As shown in Fig.
+3, directly using sparse point cloud without any point densification only represents the objects with a small number of points or fails to model the tiny objects.
+Meanwhile, in case extra points are added to the point cloud, points successfully represent the objects densely.
+Also, the quantitative result is presented in Tab.
+11.
+It shows assigning valid color features to the additional points is important to deblur the scene and reconstruct the fine details.
 Table 11: Ablation study on adding the extra points.
@@
 Methods PSNR $\uparrow$ SSIM $\uparrow$
-w/o Extra Points 22.56 0.7016
-w/ Random Colors 22.90 0.7089
-w/ Extra Points 23.80 0.7430
-Table 5: Ablation study on depth-depending pruning. Naive pruning stands for using naive points pruning from 3D-GS and Depth-based pruning stands for applying our depth-based pruning.
-Methods PSNR $\uparrow$ SSIM $\uparrow$
-Naive pruning 23.33 0.7296
-Depth-based pruning 23.80 0.7430
-5 Limitations & Future Works
-NeRF-based deblurring methods [22, 20, 28], which are developed under the assumption of volumetric rendering, are not straightforwardly applicable to rasterization-based 3D-GS [15].
-However, they can be compatible to rasterization by optimizing their MLP to deform kernels in the space of the rasterized image instead of letting MLP deform the rays and kernels in the world space.
-Although it is an interesting direction, it will incur additional costs for interpolating pixels and just implicitly transform the geometry of 3D Gaussians.
-Therefore, we believe that it will not be an optimal way to model scene blurriness using 3D-GS [15].
-Another approach to deblur the defocus blur using 3D-GS [15] can be convolving the rasterized image with conventional grid blur kernels such as gaussian blur kernel [14].
-With this approach however, the kernel expressivity is limited compared to the learnable kernels since the Gaussian blur kernel is a simple unimodal Gaussian distribution along each axis and is not learnable.
-Instead of a fixed Gaussian blur kernel, adopting a learnable grid blur kernel can also be another promising direction to address defocus blur.
-Nevertheless, as blurriness naturally varies spatially across the scene and each pixel, it is ideal to train blur kernels for every pixel from all training views.
-Thus, such method can incur high memory requirements, and its usage is limited if the rendering image is high resolution or lots of training images are provided.
-6 Conclusion
-We proposed Deblurring 3D-GS, the first defocus deblurring algorithm for 3D-GS.
-We adopted a small MLP that transforms the 3D Gaussians to model the scene blurriness.
-We also further facilitate deblurring by densifying sparse point clouds from blurry input images through additional point allocation, which uniformly distributes points in the scene and assigns color features using the K-Nearest-Neighbor algorithm.
-Also, as we applied depth-based pruning instead of naive pruning, which 3D-GS has adopted, we could preserve more points at the far edge of the scene where SfM usually struggles to extract features and fails to generate enough points.
-Through extensive experiments, we validated that our method can deblur the defocus blur while still enjoying the real-time rendering with FPS $>$ 200.
-This is because we use the MLP only during the training time, and the MLP is not involved in the inference stage, keeping the inference stage identical to the 3D-GS.
-Our method achieved state-of-the-art performance or performed comparably with current cutting-edge models evaluated under different metrics.
+w/o Extra Points 23.34 0.7374
+w/ Random Colors 23.46 0.7411
+w/ Extra Points 23.73 0.7474
+B.0.5 Training time comparison
+Tab.
+12 shows training time and the number of the Gaussians and Tab.
+13 describe the running time MLP and rasterization.
+Although MLP introduces additional costs, our approach is more efficient than prior works that generate kernels and apply convolutions.
+3D-GS hardly models the camera motion blur thus fewer 3D Gaussians are involved, dropping the training time.
+Also, the proposed method uses $M=5$, a relatively smaller number of frames compared to the existing works, leading to faster training time.
+Table 12: Comparison on training time (RTX 4090 GPU).
+Camera Motion Blur Defocus Blur
+Time (min) # Gaussians Time (min) # Gaussians
+3D-GS $\approx$ 2.1 63,549 $\approx$ 3.0 308,986
+Ours $\approx$ 13.4 139,545 $\approx$ 8.2 317,964
+Table 13: Running time of each module.
+MLP forwarding 1.33 ms Rasterization 1.00 ms
+B.0.6 Images from nearly the same viewpoints
+Two training images in Fig.
+13 are taken with nearly the same viewpoints but (a) is defocused at the near plane while (b) shows defocus blur at the far plane.
+Even if they are captured at very close views, the proposed method can deblur only the blurry regions, keeping the clean regions as they are, which highlights its view-dependent functionality.
+Figure 13: Deblurred images at nearly the same viewpoints.
+Figure 14: Qualitative results on real defocus blur dataset.
+Figure 15: Qualitative results on synthetic defocus blur dataset.
+Figure 16: Qualitative results on synthetic camera motion blur dataset.